### v1

# Intersections of matroids

Abstract.
We study simplicial complexes (hypergraphs closed under taking subsets) that are the intersection of a given number $k$ of matroids. We prove bounds on their chromatic numbers (the minimum number of edges required to cover the ground set) and their list chromatic numbers. Settling a conjecture of Király and Bérczi et. al., we prove that the list chromatic number is at most $k$ times the chromatic number. Following the footsteps of Edmonds, who considered the case $k=2$, we study three polytopes associated with $k$-tuples of matroids, and prove bounds on the distances between them. The tools used are in part topological.
1. Introduction
In a seminal paper [12] Jack Edmonds showed how combinatorial duality (min-max results) can be formulated in terms of the intersection of two matroids. The classical case is the König–Hall theorem, which can be viewed as a statement on the intersection of two partition matroids, defined on the edge-set of the graph, and whose parts are the stars in the respective sides. A matching is a set of edges belonging to the intersection of the two matroids, and a marriage is a base in one matroid that is independent in the other. Another contribution of Edmonds was setting the results in terms of polytopes associated with the matroids. The polytope $P(\mathcal{C})$ of a complex (a hypergraph closed under taking subsets) $\mathcal{C}$ is the convex hull of the characteristic functions of its edges. Edmonds showed that if $\mathcal{M}$ and $\mathcal{N}$ are matroids, then $P(\mathcal{M}\cap\mathcal{N})=P(\mathcal{M})\cap P(\mathcal{N})$.
What about the intersection of more than two matroids? These are notoriously more complex. In particular, $\min=\max$ results are no longer available. In [1, 2, 7, 19] topological tools were provided, yielding results of the form $\min\leq\max$, and sufficient conditions for matchability rather than necessary and sufficient. In this paper we continue these ideas, to study the properties of sets of matroids on the same ground set.
Particular stress will be put on the notion of coloring.
Definition 1.1.
For a complex $\mathcal{C}$, a coloring by $\mathcal{C}$ is a set of faces (edges) in $\mathcal{C}$ whose union is the ground set $V(\mathcal{C})$. The chromatic number $\chi(\mathcal{C})$ of $\mathcal{C}$ is the minimal number of faces in a coloring.
Remark 1.2.
The origin of the term “coloring” used here, is in the case that $\mathcal{C}=\mathcal{I}(G)$, the complex of independent sets in the graph $G$, for which the notion of a coloring of $\mathcal{I}(G)$, as it is defined here, coincides with the ordinary notion of proper vertex coloring of the graph $G$, commonly used in literature.
In the basic setting, we are given a set =$\mathcal{L}=\{\mathcal{M}_{1},\ldots\mathcal{M}_{k}\}$ of matroids on the same ground set $V$, and we study the relationship between certain parameters of their intersection and (the same or other) parameters of each matroid $\mathcal{M}_{i}$. These parameters are:
  1. (1)
    The chromatic number $\chi$,
  2. (2)
    $\chi^{*}$ — the fractional version of $\chi$,
  3. (3)
    $\chi_{\ell}$, the list chromatic number (to be defined below),
  4. (4)
    The matching number, which is the largest size of a set in $\bigcap_{1\leq i\leq k}\mathcal{M}_{i}$,
  5. (5)
    The fractional matching number, which is $\max\{\sum_{v\in V}f(v):f\in\bigcap_{i=1}^{k}P(\mathcal{M}_{i})\}$,
  6. (6)
    Topological connectivity (denoted by $\eta$ defined below), and
  7. (7)
    the “topological expansion number”, which is $\max_{\emptyset\neq S\subseteq V}\frac{|S|}{\eta(\mathcal{C}[S])}$ for $\mathcal{C}=\cap_{i=1}^{k}\mathcal{M}_{i}$.
We shall address the following questions:
  1. (1)
    Bounding $\chi_{\ell}(\bigcap\mathcal{L})$ in terms of $\chi(\bigcap\mathcal{L})$. We show that
    $\chi_{\ell}(\bigcap\mathcal{L})\leq k\chi(\bigcap\mathcal{L}).$
  2. (2)
    Bounding $\chi_{\ell}(\bigcap\mathcal{L})$ in terms of $\max\chi(\mathcal{M}_{i})$. We show that
    $\chi_{\ell}(\bigcap\mathcal{L})\leq(2k-1)\max_{1\leq i\leq k}\chi(\mathcal{M}_{i}).$
  3. (3)
    Bounding $\chi^{*}(\bigcap\mathcal{L})$ in terms of $\chi^{*}(\mathcal{M}_{i})$. We show that
    $\chi^{*}(\bigcap\mathcal{L})\leq k\max_{1\leq i\leq k}\chi^{*}(\mathcal{M}_{i}).$
  4. (4)
    Bounding the ratio between the polytopes $P(\bigcap\mathcal{L})$ and $\bigcap_{i=1}^{k}P(\mathcal{M}_{i})$, namely by what factor should the first be blown up to contain the second. Again, we show that a factor $k$ suffices.
The result in (1) answers the problem in [9, 21], that asked to find a constant $c$ such that $\chi_{\ell}(\bigcap\mathcal{L})\leq c\chi(\bigcap\mathcal{L})$ for every $\mathcal{L}$ consisting of two matroids.
We conjecture that in (1), (3), and (4) the factor $k$ can be replaced by $k-1$. In the case of  (3) and (4) we prove this conjecture when all $\mathcal{M}_{i}$ are partition matroids.
In the case $k=2$ the conjecture is that $\chi_{\ell}(\mathcal{C})=\chi(\mathcal{C})$ whenever $\mathcal{C}$ is the intersection of two matroids. A famous theorem of Galvin [16] is that this is true in the case of two partition matroids. The conjecture that this is true for any pair of matroids was offered in [9, 22]. The third author of the present paper proved this conjecture in the case of two generalized partition matroids [17]. As to (2), there are reasons to suspect that a factor $k-1$ suffices when complemented by an additive term.
A special case that will serve as motivation and yardstick is that in which all $\mathcal{M}_{i}$ are partition matroids. This means that for each $1\leq i\leq k$ a partition $\mathcal{P}^{i}=(P^{i}_{1},P^{i}_{2},\ldots,P^{i}_{m_{i}})$ of the ground set $V$ is given, and for $e\subseteq V$, $e\in\mathcal{M}_{i}$ if and only if $|e\cap P^{i}_{j}|\leq 1$ for every $1\leq j\leq m_{i}$. This can be described in terms of a familiar object: a $k$-partite $k$-uniform hypergraph $\mathcal{K}=\mathcal{K}(\mathcal{L})$. The vertices of $\mathcal{K}$ are the parts $P^{i}_{j}$, and each of its sides $S_{i}$ is $\{P^{i}_{j}\mid 1\leq j\leq m_{i}\}$. Every vertex $v\in V(\mathcal{L})$ corresponds to an edge $e(v)=\{P^{i}_{j}\mid 1\leq i\leq k,v\in P^{i}_{j}\}$ of $\mathcal{K}(\mathcal{L})$. Then $\bigcap\mathcal{L}=\mathcal{M}(\mathcal{K})$, the set of matchings in $\mathcal{K}$. A function $f\in\bigcap_{1\leq i\leq k}P(\mathcal{M}_{i})$ (as defined in the first paragraph) then corresponds to a fractional matching in the hypergraph $\mathcal{K}(\mathcal{L})$. Hence the name “fractional matching number” in (5) above.
In the opposite direction, given a $k$-partite hypergraph $\mathcal{H}$, we construct a collection =$\mathcal{L}=\{\mathcal{M}_{1},\dots,\mathcal{M}_{k}\}$ of $k$ partition matroids, each having  $E(\mathcal{H})$ as ground set, where each parts of $\mathcal{M}_{i}$ is a star $S_{x}$, namely the set of edges incident to a vertex $x$ in the $i$th side of $\mathcal{H}$. Such a system of partition matroids is said to be associated with $\mathcal{H}$, and is denoted by $\mathcal{L}(\mathcal{H})$. Clearly,
$\mathcal{K}(\mathcal{L}(\mathcal{H}))=\mathcal{H}.$
2. Some definitions and notation
Notation 2.1.
For a subset $A$ of $V$ let $\mathbf{1}_{A}\in\mathbb{R}^{V}$ be the characteristic function of $A$.
Notation 2.2.
For a function $f:V\rightarrow\mathbb{R}$ and $U\subseteq V$, let $f[U]=\sum_{v\in U}f(v)$.
Notation 2.3.
For functions $f,g:V\rightarrow\mathbb{R}$, $f\cdot g$ is the inner product $\sum_{v\in V}f(v)g(v)$.
Functions will also be viewed as vectors, so a real-valued function $f$ on a set $V$ is also denoted by $\vec{f}\in\mathbb{R}^{V}$.
Notation 2.4.
Given a set $A$ of vectors and a constant $\lambda$, let
$\lambda A=\{\lambda\vec{x}\mid\vec{x}\in A\}.$
Definition 2.5.
A hypergraph $\mathcal{H}$ is a collection of subsets of its vertex set (or ground set), a finite set $V=V(\mathcal{H})$. Each set $S\in\mathcal{H}$ is called an edge. If all edges are of the same size $k$ we say that $\mathcal{H}$ is $k$-uniform, or that it is a $k$-graph. For $U\subseteq V$, $\mathcal{H}[U]=\{S\in\mathcal{H}\mid S\subseteq U\}$ is the subhypergraph of $\mathcal{H}$ induced on $U$.
Definition 2.6.
A hypergraph $\mathcal{H}$ is $k$-partite if the vertex set of $\mathcal{H}$ can be divided into $k$ parts $V_{1},\dots,V_{k}$ such that for every edge $S\in\mathcal{H}$, $|S\cap V_{i}|=1$ for each $1\leq i\leq k$.
Definition 2.7.
A non-empty hypergraph $\mathcal{C}$ is called an (abstract) complex if it is closed under taking subsets, namely if $S\in\mathcal{C}$ then $T\in\mathcal{C}$ for every $T\subseteq S$. An edge of $\mathcal{C}$ is called a face.
Convention 2.8.
Unless otherwise stated, the vertex set of a complex under discussion is denoted by $V$.
Notation 2.9.
For a complex $\mathcal{C}$ and $U\subseteq V$, let $rank_{\mathcal{C}}(U)=\max_{S\in\mathcal{C}[U]}|S|$. The rank of $\mathcal{C}$, denoted by $rank(\mathcal{C})$, is $rank_{\mathcal{C}}(V)$.
A matching is a set of disjoint edges, and a cover of a hypergraph is a set of vertices meeting all edges. The set of matchings in $\mathcal{H}$ is denoted by $\mathcal{M}(\mathcal{H})$. The maximal size of a matching in $\mathcal{H}$ is denoted by $\nu(\mathcal{H})$, and the minimal size of a cover by $\tau(\mathcal{H})$. Obviously, $\tau(\mathcal{H})\geq\nu(\mathcal{H})$.
A set of vertices of $\mathcal{H}$ is said to be independent if it does not contain an edge. The complex of independent sets in $\mathcal{H}$ is denoted by $\mathcal{I}(\mathcal{H})$.
Definition 2.10.
A complex $\mathcal{M}$ is called a matroid if $\emptyset\in\mathcal{M}$, and for all $S,T\in\mathcal{M}$ satisfying $|S|<|T|$ there exists $v\in T\setminus S$ such that $S\cup\{v\}\in\mathcal{M}$.
This is equivalent to demanding that every maximal edge of $\mathcal{M}$ is of maximal size, and that this is true also for any induced subcomplex.
The edges of a matroid are said to be independent in it (a term taken from the matroid of linearly independent sets in $\mathbb{F}^{n}$ for a finite field $\mathbb{F}$, not to be confused with independent sets of vertices in a hypergraph). A base of a matroid is a maximal independent set in the matroid. A circuit $C\subseteq V$ of $\mathcal{M}$ is a minimal dependent set (the name comes from graphic matroids, namely matroids consisting of acyclic sets of edges in a graph). For $A\subseteq V(\mathcal{M})$,
$span_{\mathcal{M}}(A)=A\cup\Big{\{}x\in V\mid\{x\}\cup S\not\in\mathcal{M}\text{ for some $S\in\mathcal{M}[A]$}\Big{\}}.$
$A\subseteq V$ is called spanning if $span_{\mathcal{M}}(A)=V$.
Definition 2.11.
Given a partition of a set $V$ into parts $P_{1},\ldots,P_{m}$ and numbers $p_{1},\ldots,p_{m}$, the hypergraph $\{e\subseteq V:|e\cap P_{i}|\leq p_{i}\text{ for all }1\leq i\leq m\}$ is a matroid, called a generalized partition matroid with constraint parameters $p_{i}$. If all $p_{i}$s are $1$, it is plainly called a partition matroid.
3. Expressing complexes as intersections of matroids
Observation 3.1.
Every complex is the intersection of generalized partition matroids.
Proof.
For a subset $U$ of $V$, the complex $NC(U):=\{W\subseteq V:W\not\supseteq U\}$ is a generalized partition matroid, with two parts, $U$ and $V\setminus U$, and constraints $p_{1}=|U|-1,p_{2}=|V\setminus U|$. Noting that $\mathcal{C}=\bigcap_{U\not\in\mathcal{C}}NC(U)$, the observation follows upon. ∎
For a complex $\mathcal{C}$, let $matdim(\mathcal{C})$ be the minimum number of matroids whose intersection is $\mathcal{C}$. This is a measure (one among many) of the complexity of $\mathcal{C}$ Let $md(n)$ be the maximum of $matdim(\mathcal{C})$ over all complexes $\mathcal{C}$ on $n$ vertices.
For a hypergraph $\mathcal{H}$ let $\min(\mathcal{H})$ be the set of minimal edges in $\mathcal{H}$. Let $\chi^{\prime}(\mathcal{H})$ be the edge chromatic number of $\mathcal{H}$, i.e., $\chi^{\prime}(\mathcal{H})=\chi(\mathcal{M}({\mathcal{H}}))$ (Recall that $\mathcal{M}(\mathcal{H})$ is the set of matchings in $\mathcal{H}$).
Observation 3.2.
$matdim(\mathcal{C})\leq\chi^{\prime}(\min[\mathcal{C}^{c}])$, where $\mathcal{C}^{c}:=\{U\subseteq V(\mathcal{C})\mid U\not\in\mathcal{C}\}$. In particular, $md(n)\leq 2^{n}$.
Proof.
Let $\chi^{\prime}(\min[\mathcal{C}^{c}])=m$ and let $N_{1},\ldots,N_{m}$ be matchings whose union is the hypergraph $\min[\mathcal{C}^{c}]$. For each $1\leq i\leq m$ let $\mathcal{M}_{i}=\bigcap_{e\in N_{i}}NC(e)$. Then each $\mathcal{M}_{i}$ is a generalized partition matroid whose parts are the edges of the matching $N_{i}$, together with one part $Q=V\setminus\bigcup_{e\in N_{i}}e$, the constraint parameter of part $e$ being $|e|-1$ and that of $Q$ being $|Q|$.
We claim that $\mathcal{C}=\bigcap_{i=1}^{m}\mathcal{M}_{i}$. To see that $\mathcal{C}\subseteq\bigcap_{i=1}^{m}\mathcal{M}_{i}$, note that for every $f\in\mathcal{C}$, there does not exist $e\in N_{i}$ for some $i$ such that $e\subseteq f$, otherwise $e\in\mathcal{C}$ by the closed down property of $\mathcal{C}$, contradictory to $e\in\mathcal{C}^{c}$. Therefore $f\in\cap_{i=1}^{m}\mathcal{M}_{i}=\cap_{i=1}^{m}\cap_{e\in N_{i}}NC(e)$. To see that $\mathcal{C}\supseteq\bigcap_{i=1}^{m}\mathcal{M}_{i}$, suppose there exists $f\in\cap_{i=1}^{m}\mathcal{M}_{i}$ but $f\not\in\mathcal{C}$, then there exists $f^{\prime}\subseteq f$ such that $f^{\prime}\in\min[\mathcal{C}^{c}]$. Therefore $f^{\prime}\in N_{i}$ for some $i$, which implies $f\not\in NC(f^{\prime})$. This implies $f\not\in\mathcal{M}_{i}$, a contradiction. ∎
Notation 3.3.
Let $\mathcal{D}^{k}$ be the collection of sets =$\mathcal{L}=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{k}\}$, of matroids sharing the same ground set. For such $\mathcal{L}$, let $\bigcap\mathcal{L}=\bigcap_{i=1}^{k}\mathcal{M}_{i}$. Let $MINT_{k}=\{\bigcap\mathcal{L}\mid\mathcal{L}\in\mathcal{D}^{k}\}$.
Definition 3.4.
A complex $\mathcal{C}$ is called a flag complex if it is $2$-determined, meaning that $e\in\mathcal{C}$ if and only if $\binom{e}{2}\subseteq\mathcal{C}$.
The complex $\mathcal{I}(G)$ is a flag complex for every graph $G$. Conversely, if $\mathcal{C}$ is a flag complex, let $G(\mathcal{C})$ be the graph with vertex set $V(\mathcal{C})$ and edge set consisting of the pairs of vertices not belonging to $\mathcal{C}$. Then $\mathcal{C}=\mathcal{I}(G)$. Furthermore, for a flag complex $\mathcal{C}$, $\min[\mathcal{C}^{c}]=G(\mathcal{C})$. By Observation 3.2, if $\mathcal{C}$ is a flag complex then $matdim(\mathcal{C})\leq\chi^{\prime}(G(\mathcal{C}))$. By Vizing’s theorem, this implies that $\mathcal{C}\in MINT_{|V(\mathcal{C})|}$.
For a set $S$, let $2^{S}:=\{T\mid T\subseteq S\}$.
Example 3.5.
Let $A,B$ be disjoint sets, where $m=|B|\geq|A|$, and let $\mathcal{C}=2^{A}\cup 2^{B}$. Then $\mathcal{C}=\mathcal{I}(G)$ is the complete bipartite graph $G$ with sides $A$ and $B$, thus $\mathcal{C}$ is a flag complex. We claim that $matdim(\mathcal{C})=m$. Clearly (for example by König’s edge coloring theorem), $\chi^{\prime}(G)=\Delta(G)=|B|=m$, so $matdim(\mathcal{C})\leq m$ follows from the above observation.
To prove that $matdim(\mathcal{C})\geq m$, assume that $\mathcal{C}=\bigcap_{1\leq i\leq k}\mathcal{M}_{i}$. Fix some $a\in A$, for every $b_{i}\in B=\{b_{1},\dots,b_{m}\}$ there exists $j=j(i)\in[k]$ such that the pair $\{a,b_{i}\}$ does not belong to $\mathcal{M}_{i}$, since $\{a,b_{i}\}\not\in\mathcal{C}$. Note that if $j(i)=j(i^{\prime})=:j$ then $i=i^{\prime}$, since otherwise the matroid $\mathcal{M}_{j}$ would contain the edge $\{b_{i},b_{i^{\prime}}\}$, and none of the edges $\{a,b_{i}\},\{a,b_{i^{\prime}}\}$, which would mean that no vertex of $\{b_{i},b_{i^{\prime}}\}$ can be added to the singleton $\{a\}$ while remaining inside $\mathcal{M}_{j}$, contradicting the assumption that $\mathcal{M}_{j}$ is a matroid. We have thus shown that $k\geq m$.
Remark 3.6.
In the examples above the matroids used in the intersection are generalized partition matroids. Not every complex $\mathcal{C}$ is the intersection of $matdim(\mathcal{C})$ generalized partition matroids. An example for this is the set of the edge sets of the four stars of size 3 in $K_{4}$ on $E(K_{4})$ and their subsets. It is the intersection of the graphic matroid and the partition matroid in which the three parts are matching of size $2$. Yet it is not the intersection of two generalized partition matroid, because every such matroid on $E(K_{4})$ has $\chi\leq 2$, while by [17, Theorem 1.5] would imply $\chi(\mathcal{C})\leq 2$, which is not the case.
Observation 3.7.
$\mathcal{C}$ is a flag complex if and only if it is the intersection of partition matroids.
Proof.
Clearly, the intersection of flag complexes is a flag complex. Also obvious is that a partition matroid is a flag complex. Hence the intersection of partition matroids is a flag complex. For the other direction, if $\mathcal{C}=\mathcal{I}(G)$ for a graph $G$ then by the above $\mathcal{C}=\bigcap_{e\in G}NC(e)$ and $NC(e)$ is a partition matroid, one of its parts being $e$ and the other parts being the singleton sets $\{x\}$ for each $x\in V\setminus e$. ∎
Theorem 3.8.
$md(n)\geq\binom{n-1}{\lfloor n/2\rfloor}$.
Proof.
Let the ground set be $[n]$. Let
$\mathcal{C}=\{e\subseteq[n]:|e|\leq\lfloor n/2\rfloor\text{ or }n\not\in e\}.$
We claim that if $\mathcal{C}=\bigcap_{1\leq i\leq t}\mathcal{M}_{i}$, where $\mathcal{M}_{i}$ are matroids, then $t\geq\binom{n-1}{\lfloor n/2\rfloor}$. For every set $A\subseteq[n-1]$ of size $\lfloor n/2\rfloor$, since $A\cup\{n\}\not\in\mathcal{C}$, there exists $i=i(A)$ for which $A\cup\{n\}\not\in\mathcal{M}_{i}$. Clearly, $A$ is minimal non-$\mathcal{M}_{i}$ set, namely a circuit. Suppose there exists distinct $A_{1},A_{2}\in\binom{[n-1]}{\lfloor n/2\rfloor}$ such that $i(A_{1})=i(A_{2})=:i$, then $n$ is in the intersection of the two distinct circuits $A_{1}\cup\{n\}$ and $A_{2}\cup\{n\}$ in $\mathcal{M}_{i}$. By the circuit elimination axiom of matroids, there exists a circuit of $\mathcal{M}_{i}$ contained in $A_{1}\cup A_{2}$, which contradicts the fact that $[n-1]\in\mathcal{C}\subseteq\mathcal{M}_{i}$. Therefore the map $A\mapsto i(A)$ is injective, meaning that $t\geq\binom{n-1}{\lfloor n/2\rfloor}$. ∎
4. Topological Hall
Remark 4.1.
Below, $S^{-1}$ is interpreted as $\emptyset$ and $B^{0}$ as a single point
Definition 4.2.
A topological space $X$ is homotopically $k$-connected if for every $-1\leq i\leq k$, every continuous function from the $i$-dimensional sphere $S^{i}$ to $X$ can be extended to a continuous function from the ball $B^{i+1}$ to $X$. If $X=\emptyset$, the it is homologically $(-1)$-connected. If $X\neq\emptyset$, it is homologically $k$-connected if $\tilde{H}_{i}(X)=0$ for all $0\leq i\leq k$, where $\tilde{H}_{i}(X)$ is the reduced $i$th homology group of $X$.
Notation 4.3.
Let $\eta(X)$ (resp. $\eta_{H}(X)$) be the maximal $k$ for which $X$ is $k$-connected (resp. homologically $k$ connected), plus 2.
Theorem 4.4 (Hurwitz, [18]).
$\eta_{H}\geq\eta$. If $\eta(X)\geq 3$ (which signifies“$X$ is simply connected”), then $\eta(X)=\eta_{H}(X)$.
An abstract complex $\mathcal{C}$ has a geometric realization $||\mathcal{C}||$, obtained by positioning its vertices in general position in $\mathbb{R}^{n}$ for large enough $n$ ($n=2\cdot rank(\mathcal{C})-1$ suffices, for example any rank-2 complex, namely a graph, can be realized without fortuitous intersections in $\mathbb{R}^{3}$) and representing every face by the convex hull of its vertices, the “general position” condition guaranteeing that there are no unwanted intersections. We write $\eta(\mathcal{C})$ for $\eta(||C||)$, and $\eta_{H}(\mathcal{C})$ for $\eta_{H}(||C||)$.
Intuitively, $\eta(X)$ is the minimal dimension of a hole in $X$. For example, $\eta(S^{n})=n+1$ because the hole, namely the non-filled ball, is of dimension $n+1$. For the disk $B^{n}$, $\eta(B^{n})=\infty$. $\eta(X)=0$ means $X=\emptyset$, and $\eta(X)=1$ means that $X$ is non-empty and is not path-connected. For another example, $\eta(\mathcal{C})\geq 2$ means path-connectedness. The “2” is the size of the simplices that are used for filling those holes that are fillable. This is valid also for higher numbers, and this significance of $\eta$ makes its use natural in combinatorial settings.
Definition 4.5.
The join $\mathcal{C}*\mathcal{D}$ of two complexes on disjoint ground sets is defined as $\{A\cup B\mid A\in\mathcal{C},B\in\mathcal{D}\}$. If $V(\mathcal{C})\cap V(\mathcal{D})\neq\emptyset$, we define $\mathcal{C}*\mathcal{D}$ by first making a copy of $\mathcal{D}$ on a ground set that is disjoint from that of $\mathcal{C}$, and then taking the join of $\mathcal{C}$ and the copy of $\mathcal{D}$.
Proposition 4.6.
$\eta_{H}(\mathcal{C}*\mathcal{D})=\eta_{H}(\mathcal{C})+\eta_{H}(\mathcal{D})$.
Homotopic $\eta$ does not yield equality here, but we have the useful
Proposition 4.7.
$\eta(\mathcal{C}*\mathcal{D})\geq\eta(\mathcal{C})+\eta(\mathcal{D})$.
A basic fact, proved by Whitney conecrens the connectivity of matroids.
Proposition 4.8.
If $\mathcal{M}$ is a matroid then $\eta(\mathcal{M})=rank(\mathcal{M})$, unless $\mathcal{M}$ has a co-loop, namely an element belonging to all bases, in which case $\eta(\mathcal{M})=\infty$.
A useful generalization was proved in [2]:
Proposition 4.9.
If $\mathcal{C}\in MINT_{k}$ then $\eta(\mathcal{C})\geq\frac{rank(\mathcal{C})}{k}$.
This is sharp: let $\mathcal{K}$ be a set of $k$ disjoint edges of size $k$ each, together with an edge of size $k$ meeting all. Regarding $\mathcal{K}$ as a $k$-partite hypergraph, let $\mathcal{L}(\mathcal{K})$ be the set of $k$ partition matroids associated with $\mathcal{K}$ so that $\mathcal{M}(\mathcal{K})=\bigcap\mathcal{L}\in MINT_{k}$. Thus $\eta(\mathcal{M}(\mathcal{K}))=1$ and $rank(\mathcal{M}(\mathcal{K}))=k$.
For a general $k$-uniform $k$-partite hypergraph $\mathcal{H}$, the above argument and the proposition yield $\eta(\mathcal{M}(\mathcal{H}))\geq\nu(\mathcal{H})/k$ (This is true for every $k$-uniform hypergraph, but it will not be used here). A more general result is the following, where $\nu^{*}(\mathcal{H})$ is the fractional matching number of the hypergraph $\mathcal{H}$ (see Definition 14.2).
Theorem 4.10 (Aharoni–Berger–Meshulam [4]).
For a $k$-uniform hypergraph $\mathcal{H}$,
$\eta(\mathcal{M}(\mathcal{H}))\geq\frac{\nu^{*}(\mathcal{H})}{k}.$
We shall comment on this useful fact in Appendix B.
4.1. A topological marriage theorem
The following theorem, a topological generalization of Hall’s marriage theorem, was proved implicitly in [7] and formulated explicitly by the first author (see remark following Theorem 1.3 in [24]).
Theorem 4.11.
Let $\mathcal{C}$ be a complex, and let $V_{i}$ for $1\leq i\leq m$ be subsets of $V[\mathcal{C}]$. If $\eta(\mathcal{C}[\bigcup_{i\in I}V_{i}])\geq|I|$ for every $I\subseteq[m]$ then there exists a choice function whose image belongs to $\mathcal{C}$, namely a function $\phi:[m]\rightarrow\bigcup_{i=1}^{m}V_{i}$, such that $\phi(i)\in V_{i}$ for each $1\leq i\leq m$ and $Im(\phi)\in\mathcal{C}$.
The special case when $\mathcal{C}$ is a matroid is the famous theorem of Rado, in which $\eta$ is replaced by rank.
Meshulam [24] proved the homological version of this theorem, which by Theorem 4.4 is slightly stronger than Theorem 4.11.
Two matroidal extensions were found (independently) after the publication of [7], proving essentially the same result.
Given a matroid $\mathcal{M}$ and $X\subseteq V(\mathcal{M})$, .$\mathcal{M}.X$ is the contraction of $\mathcal{M}$ of $X$, defined as $\{\tau\subseteq X\mid\tau\cup\sigma\in\mathcal{M}\text{ for every $\sigma\in\mathcal{M}[V\setminus X]$}\}$. It is itself a matroid [26].
Theorem 4.12 (Theorem 4.5 in [2]).
Let $\mathcal{M}$ be a matroid and $\mathcal{C}$ be a complex on the same ground set $V$. If $\eta(\mathcal{C}[X])\geq rank(\mathcal{M}.X)$ for all $X\subseteq V$, then there is a base of $\mathcal{M}$ that is in $\mathcal{C}$.
Theorem 4.13.
(Kalai–Meshulam [19]) If $\eta(\mathcal{C}[S])\geq rank(\mathcal{M}[S^{c}])$ for every $S\subseteq V(\mathcal{C})$ then $\mathcal{M}\setminus\mathcal{C}\neq\emptyset$.
Theorem 4.12 provides a sufficient condition for the existence of a set that is large in $\mathcal{M}$, where “large” means here “spanning”, and small in $\mathcal{C}$, where “small” in a complex just means “belonging to it”. Theorem 4.13 provides a sufficient condition for the existence of a set that is small in $\mathcal{M}$ (belonging to it) and large in $\mathcal{C}$ (not belonging to it). The two theorems are transformable to each other, via an operation called “Alexander duality”. The Alexander dual $\mathcal{D}(\mathcal{C})$ of a complex $\mathcal{C}$ is $\{V\setminus A\mid A\notin\mathcal{C}\}$. The two results search for the same type of object, since $A\in spanning(\mathcal{M})\cap\mathcal{C}$ if and only if $V\setminus A\in\mathcal{M}^{*}\setminus\mathcal{D}(\mathcal{C})$, where $\mathcal{M}^{*}$ is the dual of $\mathcal{M}$ and $spanning(\mathcal{M})$ is the set of spanning sets in $\mathcal{M}$.
5. Expansion and colorability
As is customary, we set $\lceil\frac{0}{\infty}\rceil=0$, for $c>0$ $\lceil\frac{c}{\infty}\rceil=1$, and $\frac{c}{0}=\infty$.
Definition 5.1.
The simplicial expansion number $\Delta_{r}(\mathcal{C})$ of a complex $\mathcal{C}$ is
$\max_{\emptyset\neq S\subseteq V}\frac{|S|}{rank_{\mathcal{C}}(S)}.$
Since covering a set $S$ by edges of size at most $rank_{\mathcal{C}}(S)$ requires at least $\frac{|S|}{rank_{\mathcal{C}}(S)}$ edges,
(1) $\chi(\mathcal{C})\geq\Delta_{r}(\mathcal{C}).$
A topological version of $\Delta_{r}$ is the following.
Definition 5.2.
The topological expansion number $\Delta_{\eta}(\mathcal{C})$ of a complex $\mathcal{C}$ is
$\max_{\emptyset\neq S\subseteq V}\frac{|S|}{\eta(\mathcal{C}[S])}.$
Notation 5.3.
For a complex $\mathcal{C}$, we define $\bar{\eta}(\mathcal{C})=\min(\eta(\mathcal{C}),rank(\mathcal{C}))$.
It is sometimes helpful to replace $\eta$ by $\bar{\eta}$, which is always finite:
Definition 5.4.
The expansion number $\Delta(\mathcal{C})$ of a complex $\mathcal{C}$ is
$\max_{\emptyset\neq S\subseteq V}\frac{|S|}{\bar{\eta}(\mathcal{C}(S))}.$
Then
(2) $\Delta_{r}(\mathcal{C})\leq\Delta(\mathcal{C})\quad\text{ and }\quad\Delta_{\eta}(\mathcal{C})\leq\Delta(\mathcal{C}).$
The following was first proved by Edmonds [11], extending a theorem of Nash-Williams [25] on graph arboricity.
Theorem 5.5.
In any matroid $\mathcal{M}$ we have $\chi(\mathcal{M})=\lceil\Delta(\mathcal{M})\rceil$.
For general complexes inequality holds:
Theorem 5.6 (Corollary 8.6 in [2]).
$\chi(\mathcal{C})\leq\lceil\Delta(\mathcal{C})\rceil$ for any complex $\mathcal{C}$.
6. List coloring
A notion more general than coloring is that of list coloring. Let $\mathcal{C}$ be a complex on $V$, and let $L_{v}$ be a set of permissible colors at every $v\in V$. A list coloring with respect to the lists $L_{v}$ is a function $f:V\rightarrow\cup_{v\in V}L_{v}$ satisfying $f(v)\in L_{v}$ for every $v\in V$. It is said to be $\mathcal{C}$-respecting if $f^{-1}(c)\in\mathcal{C}$ for every color $c\in\cup_{v\in V}L_{v}$.
Definition 6.1.
The list chromatic number (also called “choosability number”) $\chi_{\ell}(\mathcal{C})$ is the minimal number $p$ such that any system of lists $(L_{v}:v\in V(\mathcal{C}))$, each of size $p$, has a $\mathcal{C}$-respecting list coloring.
For $\mathcal{C}=\mathcal{I}(G)$, the independence complex of a graph $G$, these are the familiar notions of list colorability and list chromatic number of graphs.
The system of lists $L_{v}=[p]$ for every $v\in V$ has a $\mathcal{C}$-respecting list coloring if and only if $\chi(\mathcal{C})\leq p$. Hence $\chi_{\ell}(\mathcal{C})\geq\chi(\mathcal{C})$.
It is known that the ratio $\chi_{\ell}(\mathcal{C})/\chi(\mathcal{C})$ can be arbitrarily large — for $\mathcal{C}=\mathcal{I}(K_{n,n})$ we have $\chi(\mathcal{C})=2$ and $\chi_{\ell}(\mathcal{C})=\Omega(\log n)$ [13].
Virtually the same proof as that of Theorem 5.6 yields the stronger:
Theorem 6.2.
For any complex $\mathcal{C}$,
$\chi_{\ell}(\mathcal{C})\leq\lceil\Delta_{\eta}(\mathcal{C})\rceil.$
Proof.
For any system of lists $(L_{v}:v\in V(\mathcal{C}))$ satisfying $|L_{v}|=\lceil\Delta_{\eta}(\mathcal{C})\rceil$, we shall find a $\mathcal{C}$-respecting list coloring. Let $J=\bigcup_{v\in V}L_{v}$ be the set of all the colors. For every color $j\in J$, let $F_{j}=\{v\in V(\mathcal{C})\mid j\in L_{v}\}$ be the set of all elements such that the color $j$ is in their lists. Form $W=\bigcup_{j\in J}\{j\}\times F_{j}$, which is the set of $|L_{v}|$ copies of each $v\in V(\mathcal{C})$. For each $j\in J$ let $\mathcal{C}_{j}$ be the complex $\{\{j\}\times\sigma\mid\sigma\in\mathcal{C}[F_{j}]\}$, namely a copy of $\mathcal{C}[F_{j}]$. Let $\mathcal{D}=*_{j\in J}\mathcal{C}_{j}$, namely the join of all $\mathcal{C}_{j}$ for $j\in J$, which is a complex on $W$.
For $v\in V(\mathcal{C})$, let $W_{v}=\{(j,v)\mid j\in L_{v}\}$ be the set of all the copies of $v$ in $W$. For $I\subseteq V$, we claim that
(3) $\eta(\mathcal{D}[\cup_{v\in I}W_{v}])\geq|I|.$
By the super-additivity (Proposition 4.7) of $\eta$ with respect to the join operation we have
$\eta(\mathcal{D}[\cup_{v\in I}W_{v}])\geq\sum_{j\in J}\eta(\mathcal{C}[F_{j}\cap I]).$
By the definition of $\Delta_{\eta}(\mathcal{C})$, we have
$\eta(\mathcal{C}[F_{j}\cap I])\geq\frac{|F_{j}\cap I|}{\lceil\Delta_{\eta}(\mathcal{C})\rceil}.$
Every vertex $v$ of $I$ appears in $|L_{v}|=\lceil\Delta_{\eta}(\mathcal{C})\rceil$ many $F_{j}$, hence $\sum_{j\in J}|F_{j}\cap I|=\lceil\Delta_{\eta}(\mathcal{C})\rceil\cdot|I|$. Hence
$\eta(\mathcal{D}[\cup_{v\in I}W_{v}])\geq\sum_{j\in J}\eta(\mathcal{C}[F_{j}\cap I])\geq\frac{\sum_{j\in J}|F_{j}\cap I|}{\lceil\Delta_{\eta}(\mathcal{C})\rceil}=\frac{\lceil\Delta_{\eta}(\mathcal{C})\rceil\cdot|I|}{\lceil\Delta_{\eta}(\mathcal{C})\rceil}=|I|,$
completing the proof of (3).
By Topological Hall (Theorem 4.11), there exists a choice function $\phi:V\rightarrow\cup_{v\in V}W_{v}$ such that $\phi(v)\in W_{v}$ for each $v\in V$ and $Im(\phi)\in\mathcal{D}$. Especially, $Im(\phi)$ is of the form $\bigsqcup_{j\in J}\{j\}\times\sigma_{j}$ for $\sigma_{j}\in\mathcal{C}[F_{j}]\subseteq\mathcal{C}$. Coloring every vertex $v\in V$ by the color $j=\phi(v)$ (so that $v\in\sigma_{j}$) produces then the desired $\mathcal{C}$-respecting list coloring. ∎
Together with Theorem 5.5 this yields
Theorem 6.3.
For a matroid $\mathcal{M}$, $\chi_{\ell}(\mathcal{M})=\chi(\mathcal{M})$.
This was first proved by Seymour ([27]). Here is a third proof, on top of the topological one and Seymour’s, yielding a slight strengthening:
Theorem 6.4.
Let $\mathcal{M}$ be a matroid on $V$, and let $(L_{v}:v\in V)$ be a system of lists of size $k$, of permissible colors at each vertex $v$. Let $F_{c}=\{v\mid c\in L_{v}\}$. If $\chi(\mathcal{M}[F_{c}])\leq k$ for every color $c\in\cup_{v\in V}L_{v}$ then there exists an $\mathcal{M}$-respecting list coloring of $\mathcal{M}$.
Proof.
Let $C=\cup_{v\in V}L_{v}$ be the set of all colors. Consider a bipartite graph $\Gamma$ whose sides are $V$ and $C$, and whose edge set is $\{vc\mid v\in V,c\in L_{v}\}$. Let $\mathcal{P}$ be the partition matroid on $E(\Gamma)$ whose parts are the stars $S_{v}=\{vc\mid c\in L_{v}\}$ for $v\in V$. Let $\mathcal{Q}$ be the matroid consisting of those sets $F$ of edges in $\Gamma$ satisfying the condition that for every color $c\in C$ there holds
$\{v\in V\mid vc\in F\}\in\mathcal{M}.$
Note that $\mathcal{Q}$ is isomorphic to $*_{c\in C}\mathcal{M}$. Then for each $K\subseteq E(\Gamma)$ we have
(4) $rank_{\mathcal{Q}}(K)=\sum_{c\in C}rank_{\mathcal{M}}(\bigcup K\cap F_{c}).$
We need to show that $rank(\mathcal{P}\cap\mathcal{Q})=|V|$. Indeed, if $Z\in\mathcal{P}\cap\mathcal{Q}$ is of size $|V|$, then coloring each vertex $v$ by the color $c$ for which $vc\in Z$ (which is well-defined since $Z\in\mathcal{P}$ and $|Z|=|V|$) yields the desired $\mathcal{M}$-respecting list coloring (by the property $Z\in\mathcal{Q}$).
By Edmonds’ two-matroids intersection theorem, this is equivalent to the condition that for any partition of $E(\Gamma)$ into two sets $X,Y$ there holds
(5) $rank_{\mathcal{P}}(X)+rank_{\mathcal{Q}}(Y)\geq|V|.$
Claim 6.5.
The minimum of LHS of (5) can be attained by $\bar{X}=\cup_{v\in V_{1}}S_{v}$ for some $V_{1}\subseteq V$.
Proof of the claim.
Assume the minimum of LHS of (5) is obtained by some $X\subseteq E(\Gamma)$ and $Y=E(\Gamma)\setminus X$. We define $T=\cup X\cap V$, $\bar{X}=\cup_{v\in T}S_{v}$, and $\bar{Y}=E(\Gamma)\setminus\bar{X}$. On the one hand, $rank_{\mathcal{P}}(\bar{X})=rank_{\mathcal{P}}(X)=|T|$. On the other hand, since $\bar{Y}\subseteq Y$, by (4) we have
$rank_{\mathcal{Q}}(\bar{Y})=\sum_{c\in C}rank_{\mathcal{M}}(\cup\bar{Y}\cap F_{c})\leq\sum_{c\in C}rank_{\mathcal{M}}(\cup Y\cap F_{c})=rank_{\mathcal{Q}}(Y).$
Therefore
$rank_{\mathcal{P}}(\bar{X})+rank_{\mathcal{Q}}(\bar{Y})\leq rank_{\mathcal{P}}(X)+rank_{\mathcal{Q}}(Y).$
By the minimality property of $X$, setting $V_{1}=T$ proves the claim. ∎
By the above claim, we may assume $X=\cup_{v\in T}S_{v}$ for some $T\subseteq V$. Then $\cup Y\cap F_{c}=F_{c}\setminus T$ for each $c\in C$. By (4) this implies that (5) is equivalent to the condition that for every $T\subseteq V$ there holds
(6) $|T|+\sum_{c\in C}rank_{\mathcal{M}}(F_{c}\setminus T)\geq|V|.$
Write $A=V\setminus T$. The condition is that $\sum_{c\in C}rank_{\mathcal{M}}(F_{c}\cap A)\geq|A|$. We have to show this for every $A\subseteq V$, but without loss of generality, we may assume that $A=V$, because in the following proof, the general case can be obtained by modifying it and looking at $\mathcal{M}[A]$ this is allowed since $\chi(\mathcal{M}[A])\leq\chi(\mathcal{M})$.
So, we have to show that
(7) $\sum_{c\in C}rank_{\mathcal{M}}(F_{c})\geq|V|.$
The assumption $|L_{v}|=k$ means that every vertex of $V$ belongs to $k$ sets $F_{c}$. Hence, together with the fact that
$rank_{\mathcal{M}}(F_{c})\geq\frac{|F_{c}|}{\chi(\mathcal{M}[F_{c}])}\geq\frac{|F_{c}|}{k},$
changing the order of summation, we have
$\sum_{c\in C}rank_{\mathcal{M}}(F_{c})\geq\sum_{c\in C}\frac{|F_{c}|}{k}=\sum_{v\in V}\sum_{c\in C}\sum_{F_{c}\ni v}\frac{1}{k}=\sum_{v\in V}k\frac{1}{k}=|V|,$
which completes the proof. ∎
As already mentioned, Theorem 6.2 settles a conjecture in [9, 21]: it was conjectured that there exists a constant $c$ such that $\chi_{\ell}(\mathcal{C})\leq c\chi(\mathcal{C})$ for every $\mathcal{C}\in MINT_{2}$.
Theorem 6.6.
If $\mathcal{C}\in MINT_{k}$ then $\chi_{\ell}(\mathcal{C})\leq k\chi(\mathcal{C})$.
Proof.
Assume that $S\subseteq V(\mathcal{C})$ attains the maximum in the definition of $\Delta_{\eta}(\mathcal{C})$, namely $\frac{|S|}{\eta(\mathcal{C}[S])}=\Delta_{\eta}(\mathcal{C})$. By Proposition 4.9, $rank_{\mathcal{C}}(S)\leq k\eta(\mathcal{C}[S])$. Together with (1), this yields
$k\chi(\mathcal{C})\geq k\chi(\mathcal{C}[S])\geq\Big{\lceil}k\frac{|S|}{rank_{\mathcal{C}}(S)}\Big{\rceil}\geq\Big{\lceil}\frac{|S|}{\eta(\mathcal{C}[S])}\Big{\rceil}=\lceil\Delta_{\eta}(\mathcal{C})\rceil\geq\chi_{\ell}(\mathcal{C}),$
which completes the proof. ∎
Conjecture 6.7.
[9, 22] If $\mathcal{C}\in MINT_{2}$ then $\chi_{\ell}(\mathcal{C})=\chi(\mathcal{C})$.
Galvin proved the conjecture for $\mathcal{C}=\mathcal{M}\cap\mathcal{N}$ for partition matroids $\mathcal{M}$ and $\mathcal{N}$. The third author [17] recently proved the conjecture for generalized partition matroids. More cases were dealt with in in [9, 22].
7. Bounding $\Delta_{\eta}(\bigcap\mathcal{L})$ (and hence $\chi_{\ell}(\bigcap\mathcal{L})$) for the intersection of partition matroids
In this section we prove an upper bound on $\Delta_{\eta}$, thereby also on $\chi_{\ell}$, of the intersection of $k$ partition matroids. We conjecture that the result holds also for the intersection of general matroids, but in the general case we can only prove a weaker result - this will be done in the next section.
Theorem 7.1.
If =$\mathcal{L}=\{M_{1},\ldots,\mathcal{M}_{k}\}\in\mathcal{D}^{k}$ is a set of $k$ partition matroids, then
(8) $\Delta_{\eta}(\bigcap\mathcal{L})\leq k\max_{1\leq i\leq k}\Delta_{r}(\mathcal{M}_{i}).$
Proof.
Let $\mathcal{H}=\mathcal{K}(\mathcal{L})$, the $k$-partite hypergraph associated with $\mathcal{L}$ (see the definition in Section 1). Let $d=\max_{1\leq i\leq k}\Delta_{r}(\mathcal{M}_{i})$. Then the maximal degree in $\mathcal{H}$ is $d$, hence putting weight $\frac{1}{d}$ on every edge of $\mathcal{H}$ yields a fractional matching, so $\nu^{*}(\mathcal{H})\geq\frac{|E(\mathcal{H})|}{d}$. By Theorem 4.10 this implies that $\eta(\mathcal{M}(\mathcal{H}))\geq\frac{|E(\mathcal{H})|}{kd}$. Moving sides gives $\frac{|E(\mathcal{H})|}{\eta(\mathcal{M}(\mathcal{H}))}\leq kd$. By the same token, this is true for any subhypergraph of $\mathcal{H}$, yielding (8). ∎
This inequality is sharp, as shown by the following example.
Example 7.2.
The $k$-uniform affain plane is obtained by removing one line $L$ of the projective plane of uniformity $k+1$ (if such exists) and removing all vertices in the line $L$ from other lines. Thus the affine plane is a $k$-uniform hypergraph on $k^{2}$ vertices that is union of $k+1$ matchings $M_{1},\ldots,M_{k+1}$, each of size $k$, that are pairwise cross-intersecting, namely $e\cap f\neq\emptyset$ whenever $e,f$ belong to different $M_{i}$s. For example, the affine plane for $k=2$ is obtained from Fano plane in the above way, and the affine plane is a copy of $K_{4}$.
Choose one of the $k+1$ matchings, say $M_{1}$, and remove its edges (keeping the vertices). What remains is a $k$-partite hypergaph, which we name $Q_{k}$, whose sides are the edges of $M_{1}$. It is $k$-regular, namely each vertex is contained in $k$ edges. Also, $|E(Q_{k})|=k^{2}$. For example, $Q_{2}=C_{4}$.
Let =$\mathcal{L}=\{\mathcal{M}_{1},\dots,\mathcal{M}_{k}\}$ be the system of partition matroids associated with $Q_{k}$. By the pairwise cross-intersecting property, we have $\eta(\bigcap\mathcal{L})=\eta(\mathcal{M}(Q_{k}))=1$ and thus $\Delta_{\eta}(\bigcap\mathcal{L})=|Q_{k}|=k^{2}$. On the other hand, the $k$-regularity of $Q_{k}$ means $\Delta_{r}(\mathcal{M}_{i})=k$ so that $\Delta_{\eta}(\bigcap\mathcal{L})=k^{2}=k\max_{1\leq i\leq k}\Delta_{r}(\mathcal{M}_{i})$.
Combining Theorem 6.2, (2) and Theorem 5.5 yields:
Corollary 7.3.
If $\mathcal{C}$ is the intersection of $k$ partition matroids $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ on the same ground set, then $\chi_{\ell}(\mathcal{C})\leq k\max_{1\leq i\leq k}\chi(\mathcal{M}_{i})$.
8. Bounding $\Delta_{\eta}(\bigcap\mathcal{L})$ (and hence $\chi_{\ell}(\bigcap\mathcal{L})$) for the intersection of general matroids
8.1. A bound
As mentioned above, we suspect that Theorem 7.1 is also valid for $k$-tuples of general matroids, but we can only prove the weaker:
Theorem 8.1.
Let =$\mathcal{L}=\{\mathcal{M}_{1},\dots,\mathcal{M}_{k}\}\in\mathcal{D}^{k}$. Then
(9) $\Delta_{\eta}(\cap_{i=1}^{k}\mathcal{M}_{i})\leq(2k-1)\max_{1\leq i\leq k}\Delta_{r}(\mathcal{M}_{i}).$
Combining Theorem 6.2, (2), and Theorem 5.5 yields
Corollary 8.2.
Let =$\mathcal{L}=\{\mathcal{M}_{1},\dots,\mathcal{M}_{k}\}\in\mathcal{D}^{k}$. Then
$\chi_{\ell}(\cap_{i=1}^{k}\mathcal{M}_{i})\leq(2k-1)\max_{1\leq i\leq k}\chi(\mathcal{M}_{i}).$
The proof of [2, Theorem 8.9] together with Theorem 6.2 shows that:
Theorem 8.3.
For $\{\mathcal{M}_{1},\mathcal{M}_{2}\}\in\mathcal{D}^{2}$, $\chi_{\ell}(\mathcal{M}_{1}\cap\mathcal{M}_{2})\leq 2\max_{1\leq i\leq 2}\chi(\mathcal{M}_{i})$.
In [10] this was strengthened to:
$\chi_{\ell}(\mathcal{M}_{1}\cap\mathcal{M}_{2})\leq\chi(\mathcal{M}_{1})+\chi(\mathcal{M}_{2}).$
Conjecture 8.4.
For =$\mathcal{L}=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{k}\}\in\mathcal{D}^{k}$ there holds
$\chi(\bigcap\mathcal{L})\leq k\max_{1\leq i\leq k}\chi(\mathcal{M}_{i}).$
In [2], it was conjectured that: $\chi(\mathcal{M}_{1}\cap\mathcal{M}_{2})\leq\max(\chi(\mathcal{M}_{1}),\chi(\mathcal{M}_{2})+1)$, and it is tempting to extend the conjecture to:
Conjecture 8.5.
For =$\mathcal{L}=\{\mathcal{M}_{1},\ldots,\mathcal{M}_{k}\}\in\mathcal{D}^{k}$ there holds
$\chi(\bigcap\mathcal{L})\leq(k-1)\max_{1\leq i\leq k}\chi(\mathcal{M}_{i})+1.$
Let us now go back to the proof of Theorem 8.1. As a first step we prove it for partition matroids, ignoring the fact that in this case we have at hand a stronger result. The reason is that this proof contains the ideas of the general result in a transparent form.
Notation 8.6.
For a hypergraph $\mathcal{H}$ with vertex set $V$ and edge set $E$, for an edge $e$ of $\mathcal{H}$, let $\mathcal{H}-e$ be the hypergraph with vertex set $V$ and edge set $E\setminus\{e\}$. For a vertex subset $X\subseteq V$, let $\mathcal{H}/X$ be the contraction hypergraph whose vertex set is $V\setminus X$ and whose edge set is $\{f\setminus X:f\in E,f\not\subseteq X\}$.
8.2. The Meshulam game for graphs
A useful tool for obtaining lower bounds on $\eta(\mathcal{I}(G))$ for a graph $G$ is a theorem of Meshulam.
Theorem 8.7.
[24] If $e\in E(G)$ then
(10) $\eta\big{(}\mathcal{I}(G)\big{)}\geq\min\Big{(}\eta\big{(}\mathcal{I}(G-e)\big{)},\eta\big{(}\mathcal{I}(G/e)\big{)}+1\Big{)}.$
In [5, 3] this result was formulated in terms of a game between two players: CON and NON. CON (who wants to establish high connectivity) offers at each step a so-far non-removed edge $e$, and NON chooses whether to remove it or to “explode” it, namely to remove the vertices of $e$ and all the neighboring vertices of $e$. The independence complex of a graph after exploding $e$ is easily seen to be the same as that of $G/e$. The value of the game is the maximal number of explosions, assuming best strategies for both players. The theorem says that this is a lower bound on $\eta(\mathcal{I}(G))$.
For a set $Z$ of vertices in a graph $G$ let $N_{G}(Z)=\{u\in V(G)\mid\{u,z\}\in E(G)\text{ for some $z\in Z$}\}$ be the set of neighbors of $Z$ in $G$.
Definition 8.8.
A set $S$ of vertices in a graph $G=(V,E)$ is said to dominate a vertex subset $T$ if for every $t\in T$ there exists $s\in S$ such that $\{s,t\}\in E$. The set of sets dominating $T$ is denoted by $Dom(T)$. A set $\mathcal{F}$ of edges of $G$ is dominating if $\bigcup\mathcal{F}$ is dominating.
Notation 8.9.
Let $\gamma(T):=\min\{|D|:D\in Dom(T)\}$. Let $\gamma^{E}(G)$ be the minimal size of a dominating set of edges of $G$, and be $\infty$ if no such dominating set exists.
Meshulam’s theorem easily implies:
Theorem 8.10.
For a graph $G$,
(11) $\eta(\mathcal{I}(G))\geq\gamma^{E}(G).$
Proof.
Let CON offer NON edges of $G$ greedily, namely in any order. When the game ends, either the graph has an empty edge set and non-empty vertex set, in which case $\eta(\mathcal{I}(G))=\infty\geq\gamma^{E}(G)$, or the graph has an empty vertex set, meaning that all vertices have been dominated by the edges chosen so far, so $\eta(\mathcal{I}(G))$ is at least the number of explosions, which is at least $\gamma^{E}(G)$. ∎
Remark 8.11.
There is also a homotopic proof, essentially given in [6].
Definition 8.12.
For a hypergraph $\mathcal{H}$, let $L(\mathcal{H})$ be the line graph of $\mathcal{H}$, i.e., the vertex set of $L(\mathcal{H})$ is $E(\mathcal{H})$ and two vertices of $L(\mathcal{H})$ are adjacent if and only if, as edges of $\mathcal{H}$, they intersect.
Clearly, $\mathcal{M}(\mathcal{H})=\mathcal{I}(L(\mathcal{H}))$. The lower bound on $\eta(\mathcal{I}(G))$ ($G$ a graph) used in the proof above of Theorem 7.1 belongs to the family of “how hard it is to dominate an independent set” (in the case of line graphs, how hard it is to dominate matchings, or fractional matchings). Theorem 8.10, used in the proof below of Theorem 8.1, belongs to a second type - lower bounds on $\eta(\mathcal{I}(G))$ formulated in terms of domination of the entire vertex set. Proving Theorem 7.1 for general matroids may require matroidal versions of Theorem 4.10.
8.2.1. Bounding $\Delta_{\eta}(\bigcap\mathcal{L})$ for partition matroids — a second approach
When the matroids $\mathcal{M}_{i}$ are partition matroids, Theorem 8.1 boils down to (11), applied to the line graph $G$ of $\mathcal{H}=\mathcal{K}(\mathcal{L})$ (a reminder:  $\mathcal{K}(\mathcal{L})$ is the $k$-partite hypergraph associated with $\mathcal{L}$). In this case, an edge of $G$ consists of two intersecting edges of $\mathcal{H}$, each of size $k$, thus containing together at most $12k-$ 1 vertices of $\mathcal{H}$. And since the degree of each vertex of $\mathcal{H}$ is at most $t:=\max_{1\leq i\leq k}\Delta_{r}(\mathcal{M}_{i})$, every edge of $G$ dominates at most $(2k-1)t$ vertices of $G$. Thus by (11) in Theorem 8.10,
$\eta(\bigcap\mathcal{L})=\eta(\mathcal{M}(\mathcal{H}))=\eta(\mathcal{I}(G))\geq\gamma^{E}(G)\geq\frac{|V(\bigcap\mathcal{L})|}{(2k-1)t},$
and since this argument applies also to every induced subgraph of $G$, the theorem follows. Note that in the game, at each step CON offers a pair of intersecting edges in $\mathcal{H}$, which is a circuit in one of the matroids. Having this strategy in mind will make the general argument more transparent.
8.3. A hypergraph Meshulam game
To extend the proof to general matroids, we need a hypergraph version of Meshulam’s theorem, proved in [10].
Theorem 8.13.
Let $\mathcal{H}$ be a hypergraph and let $e$ be a containment-wise minimal edge of $\mathcal{H}$. Then
(12) $\eta(\mathcal{I}(\mathcal{H}))\geq\min\Big{(}\eta(\mathcal{I}(\mathcal{H}-e)),\;\eta(\mathcal{I}(\mathcal{H}/e))+|e|-1\Big{)}.$
The proof uses a Mayer–Vietoris-like inequality.
Lemma 8.14 (Lemma 2.3 in [2]).
For any pair $\mathcal{A},\mathcal{B}$ of complexes,
$\eta(\mathcal{A})\geq\min(\eta(\mathcal{A}\cup\mathcal{B}),\eta(\mathcal{A}\cap\mathcal{B})).$
The homological version of the lemma is a corollary of the Mayer–Vietoris theorem. Its homotopic version can be proved directly, as is done in [2].
Proof of Theorem 8.13. .
Let $\mathcal{A}=\mathcal{I}(\mathcal{H})$ and $\mathcal{B}=2^{e}*\mathcal{I}(\mathcal{H}/e)$. Then $\mathcal{A}\cup\mathcal{B}=\mathcal{I}(\mathcal{H}-e)$: $\mathcal{A}\subseteq\mathcal{I}(\mathcal{H}-e)$. For $S\in\mathcal{B}$, assume $S=e^{\prime}\sqcup T$, where $e^{\prime}\subseteq e$ and $T\in\mathcal{I}(\mathcal{H}/e)$. Suppose $S$ is not an independent set in $\mathcal{H}-e$, i.e., there exists $f\in\mathcal{H}-e$ such that $f\subseteq S$. Since $f\not\subseteq e$, then $f\setminus e\subseteq T$, which together with $f\setminus e\in\mathcal{H}/e$ is contradictory to the assumption that $T\in\mathcal{I}(\mathcal{H}/e)$. Therefore $S\in\mathcal{I}(\mathcal{H}-e)$ and $\mathcal{A}\cup\mathcal{B}\subseteq\mathcal{I}(\mathcal{H}-e)$.
On the other hand, for $S\in\mathcal{I}(\mathcal{H}-e)$, if $e\not\subseteq S$, then $S\in\mathcal{A}$. Otherwise $S=e\sqcup(S\setminus e)$ and $S\setminus e\in\mathcal{I}(\mathcal{H}/e)$, since for any $f\in\mathcal{H}$ with $f\neq e$, $f\not\subseteq S$ and then $f\setminus e\not\subseteq S\setminus e$. In this case $S=e\sqcup(S\setminus e)\in 2^{e}*\mathcal{I}(\mathcal{H}/e)=\mathcal{B}$. Therefore $\mathcal{I}(\mathcal{H}-e)\subseteq\mathcal{A}\cup\mathcal{B}$.
And $\mathcal{A}\cap\mathcal{B}=(2^{e}\setminus\{e\})*\mathcal{I}(\mathcal{H}/e)$: for $S\in(2^{e}\setminus\{e\})*\mathcal{I}(\mathcal{H}/e)$, assume $S=e^{\prime}\sqcup T$, where $e^{\prime}\subsetneqq e$ and $T\in\mathcal{I}(\mathcal{H}/e)$. Then obviously $S\in\mathcal{B}$. And $S\in\mathcal{A}$, since $e\not\subseteq S$ and for any other $f\in\mathcal{H}$, $(f\setminus e)\not\subseteq T$ by $T\in\mathcal{I}(\mathcal{H}/e)$. Therefore $f=(f\cap e)\sqcup(f\setminus e)\not\subseteq S$. Therefore $(2^{e}\setminus\{e\})*\mathcal{I}(\mathcal{H}/e)\subseteq\mathcal{A}\cap\mathcal{B}$.
On the other hand, for $S\in\mathcal{A}\cap\mathcal{B}$, since $S\in\mathcal{B}$, we may assume $S=e^{\prime}\sqcup S^{\prime}$, where $e^{\prime}\subseteq e$ and $S^{\prime}\in\mathcal{I}(\mathcal{H}/e)$. Since $S\in\mathcal{A}$, then $e\not\subseteq e^{\prime}$, which implies $e^{\prime}\subsetneqq e$ and then $S\in(2^{e}\setminus\{e\})*\mathcal{I}(\mathcal{H}/e)$. Therefore $\mathcal{A}\cap\mathcal{B}\subseteq(2^{e}\setminus\{e\})*\mathcal{I}(\mathcal{H}/e)$.
Applying Lemma 8.14, Proposition 4.7, and Proposition 4.8 for $\eta(2^{e}\setminus\{e\})=|e|-1$, we have
$\eta(\mathcal{I}(\mathcal{H}))$ $\geq\min\Big{(}\eta(\mathcal{I}(\mathcal{H}-e)),\;\eta\big{(}(2^{e}\setminus\{e\})*\mathcal{I}(\mathcal{H}/e)\big{)}\Big{)}$
$\geq\min\Big{(}\eta(\mathcal{I}(\mathcal{H}-e)),\;\eta(2^{e}\setminus\{e\})+\eta(\mathcal{I}(\mathcal{H}/e))\Big{)}$
$=\min\Big{(}\eta(\mathcal{I}(\mathcal{H}-e)),\;|e|-1+\eta(\mathcal{I}(\mathcal{H}/e))\Big{)},$
proving the theorem. ∎
Definition 8.15.
Given a hypergraph $\mathcal{H}$, a sequence $\mathcal{K}=(e_{1},\ldots,e_{p})$ of edges of $\mathcal{H}$ is dominating if for every vertex $v\in V\setminus\bigcup\mathcal{K}$ there exists an edge $f$ of $\mathcal{H}$ such that $f\setminus\bigcup\mathcal{K}=\{v\}$. It is called frugal if $|e_{i}\setminus\bigcup_{\ell<i}e_{\ell}|>1$ for every $1\leq i\leq p$.
Notation 8.16.
$\gamma^{E}(\mathcal{H})$ is the minimum of $|\bigcup\mathcal{K}|-|\mathcal{K}|$ over all $\mathcal{K}$ that are both frugal and dominating ($\infty$, if there is no such $\mathcal{K}$).
Theorem 8.17.
$\eta(\mathcal{I}(\mathcal{H}))\geq\gamma^{E}(\mathcal{H})$ for any hypergraph $\mathcal{H}$.
Proof.
Observe that if a hypergraph $\mathcal{G}$ has an edge of size 1, say $\{v\}$, then
(13) $\mathcal{I}(\mathcal{G})=\mathcal{I}\Big{(}\mathcal{G}[V(\mathcal{G})\setminus\{v\}]\Big{)}.$
To prove the theorem, we apply the following algorithm. Let $\mathcal{H}_{0}=\mathcal{H}$. Suppose that the hypergraph $\mathcal{H}_{j}$ is defined for some $j\geq 0$. Delete all singleton edges and the vertices forming them. By (13) this does not change the independence complex. All edges of the resulting hypergraph  $\mathcal{H}_{j}^{*}$ have sizes at least 2. If  $E(\mathcal{H}_{j}^{*})=\emptyset$ while $V(\mathcal{H}_{j}^{*})\neq\emptyset$, we stop. In this case
(14) $\eta(\mathcal{I}(\mathcal{H}_{j}^{*}))=\infty;$
Or we stop if $V(\mathcal{H}_{j}^{*})=\emptyset$, in which case
(15) $\eta(\mathcal{I}(\mathcal{H}_{j}^{*}))=0.$
If neither of these cases occurs, we proceed. Choose a containment-wise minimal edge $f_{j}$ of $\mathcal{H}_{j}^{*}$. Then $|f_{j}|\geq 2$ and by Theorem 8.13
$\eta(\mathcal{I}(\mathcal{H}_{j}^{*}))\geq\min\Big{(}\eta(\mathcal{I}(\mathcal{H}_{j}^{*}-f_{j})),\;\eta(\mathcal{I}(\mathcal{H}_{j}^{*}/f_{j}))+|f_{j}|-1\Big{)}.$
If $\eta(\mathcal{I}(\mathcal{H}_{j}^{*}))\geq\eta(\mathcal{I}(\mathcal{H}_{j}^{*}-f_{j}))$, then we set $\mathcal{H}_{j+1}=\mathcal{H}_{j}^{*}-f_{j}$; otherwise we set $\mathcal{H}_{j+1}=\mathcal{H}_{j}^{*}/f_{j}$. And we repeat this process.
Clearly we will finally stop either because the edge set becomes empty and the vertex set is not empty, or the vertex set becomes empty. If the former happens, by (14), inductively we have
$\eta(\mathcal{I}(\mathcal{H}))=\eta(\mathcal{I}(\mathcal{H}_{0}))\geq\infty\geq\gamma^{E}(\mathcal{H})$
so that the theorem is true. If the latter happens, by (15), let $e_{i}^{*}$ be those $f_{i}$ that are contracted during the algorithm and assume they are $(e_{1}^{*},\dots,e_{p}^{*})$. Then by the construction,
(16) $\eta(\mathcal{I}(\mathcal{H}))=\eta(\mathcal{I}(\mathcal{H}_{0}))\geq\sum_{i=1}^{p}(|e_{i}^{*}|-1).$
Furthermore, since the deletions of vertices or edges does not change the remaining edges, while contracting $f_{i}$ is deleting $f_{i}$ from other edges, then we know that there exists edges $e_{1},\dots,e_{p}$ of $\mathcal{H}_{0}=\mathcal{H}$ such that for each $1\leq i\leq p$
$e_{i}\setminus\cup_{\ell<i}e_{\ell}=e_{i}^{*}.$
Let $\mathcal{K}=(e_{1},\dots,e_{p})$. Since $|e_{i}^{*}|\geq 2$, we have $\mathcal{K}$ is frugal. And for every $v\in V(\mathcal{H})\setminus\bigcup\mathcal{K}$$v$ is a deleted vertex during the algorithm, which means during the algorithm there exists an edge of size one consisting of $v$. Thus for the same reason as above, there exists an edge $f$ of $\mathcal{H}$ such that $f\setminus\bigcup\mathcal{K}=\{v\}$. Therefore $\mathcal{K}$ is dominating. And we have
$|\bigcup\mathcal{K}|-|\mathcal{K}|=\sum_{i=1}^{p}|e_{i}\setminus\cup_{\ell<i}e_{\ell}|-p=\sum_{i=1}^{p}(|e_{i}^{*}|-1).$
Therefore by (16), in this case we also have
$\eta(\mathcal{I}(\mathcal{H}))\geq\gamma^{E}(\mathcal{H}),$
which completes the proof. ∎
8.4. Proof of Theorem 8.1
Observation 8.18.
For any circuit $C$ in a matroid $\mathcal{M}$, any  $v\in C$, and any subset $T$ of the ground set of $\mathcal{M}$,
$span_{\mathcal{M}}(T\cup C)=span_{\mathcal{M}}(T\cup C\setminus\{v\}).$
Proof.
$span_{\mathcal{M}}(T\cup C\setminus\{v\})\subseteq span_{\mathcal{M}}(T\cup C)$ is immediate. For the other direction, first, the element $v$ in $span_{\mathcal{M}}(T\cup C)$ is also in $span_{\mathcal{M}}(T\cup C\setminus\{v\})$, since $C\setminus\{v\}\in\mathcal{M}$ and $C=(C\setminus\{v\})\cup\{v\}\not\in\mathcal{M}$. For any $u\in span_{\mathcal{M}}(T\cup C)\setminus(T\cup C)$, there exists a base $B\in\mathcal{M}[T\cup C]$ such that $\{u\}\cup B\not\in\mathcal{M}$. Let $B^{\prime}=B\setminus C$. We claim that there exists $B^{\prime\prime}\subseteq B^{\prime}$ such that $I:=B^{\prime\prime}\cup C\setminus\{v\}\in\mathcal{M}[T\cup C\setminus\{v\}]$, $|I|=|B|$, and $\{u\}\cup I\not\in\mathcal{M}$. This implies $u\in span_{\mathcal{M}}(T\cup C\setminus\{v\})$ and completes the proof that $span_{\mathcal{M}}(T\cup C)\subseteq span_{\mathcal{M}}(T\cup C\setminus\{v\})$.
To prove the claim, we shall first construct $I$. Note that the base $B$ in $\mathcal{M}[T\cup C]$ has cardinality as least $|C\setminus\{v\}|=|C|-1$, so adding some elements of $B\setminus(C\setminus\{v\})$ one by one to $C\setminus\{v\}$ extends $C\setminus\{v\}$ to an independent set $I$, which has size $|B|$. Since $C\setminus\{v\}$ is a base in $\mathcal{M}[C]$, the elements added must come from $B^{\prime}$ and we denote that set of elements $B^{\prime\prime}$, which guarantees $I\in\mathcal{M}[T\cup C\setminus\{v\}]$. It remains to show $\{u\}\cup I\not\in\mathcal{M}$. Suppose not. Then $B$ can extend to an independent in $\mathcal{M}$ by adding one element of $\{u\}\cup I\setminus B$. That element cannot be $u$ since $\{u\}\cup B\not\in\mathcal{M}$, and cannot be any element of $I$ since $B$ is a maximal independent set in $\mathcal{M}[T\cup C]$. We get a contradiction. ∎
Definition 8.19.
The circuit hypergraph $CIRC(\mathcal{M})$ of a matroid $\mathcal{M}$ is the set of circuits in $\mathcal{M}$.
By the definition of “circuit”, if $\mathcal{H}=CIRC(\mathcal{M})$ then $\mathcal{M}=\mathcal{I}(\mathcal{H}).$
Proof of Theorem 8.1.
Let $t=\max_{1\leq i\leq k}\Delta_{r}(\mathcal{M}_{i})$, $\mathcal{C}=\cap_{i=1}^{k}\mathcal{M}_{i}$, and $V$ be the common ground set of the matroids. First we prove that
$\frac{|V|}{\eta(\mathcal{C})}\leq(2k-1)t.$
Let $\mathcal{H}_{i}=CIRC(\mathcal{M}_{i})$ for $1\leq i\leq k$ and $\mathcal{H}=\cup_{i=1}^{k}\mathcal{H}_{i}$. Then
$\mathcal{C}=\mathcal{I}(\mathcal{H}).$
We may assume that $\gamma^{E}(\mathcal{H})$ is finite, otherwise by Theorem 8.17, $\eta(\mathcal{C})=\infty$ and we have done. Let $\mathcal{K}=(e_{1},\dots,e_{p})$ be a dominating and frugal sequence of edges of $\mathcal{H}$ for which
(17) $\gamma^{E}(\mathcal{H})=|\bigcup\mathcal{K}|-|\mathcal{K}|=\sum_{j=1}^{p}|e_{j}\setminus\cup_{\ell<j}e_{\ell}|-p=\sum_{j=1}^{p}(|e_{j}\setminus\cup_{\ell<j}e_{\ell}|-1).$
Since $\mathcal{K}$ is dominating, for any $v\in V\setminus\bigcup\mathcal{K}$ there exists an edge $f\in\mathcal{H}$ such that
$f\setminus\bigcup\mathcal{K}=\{v\}.$
Let $i$ be such that $f\in\mathcal{H}_{i}=CIRC(\mathcal{M}_{i})$. Then $v\in span_{\mathcal{M}_{i}}(f\setminus\{v\})$, and then
$v\in\bigcup\mathcal{K}\cup f\subseteq span_{\mathcal{M}_{i}}(\bigcup\mathcal{K}\cup f\setminus\{v\})=span_{\mathcal{M}_{i}}(\bigcup\mathcal{K}).$
Therefore
(18) $V\subseteq\cup_{i=1}^{k}span_{\mathcal{M}_{i}}(\bigcup\mathcal{K}).$
For each $1\leq j\leq p$, noting that $e_{j}$ is an edge of $\mathcal{H}=\cup_{i=1}^{k}\mathcal{H}_{i}$, let
$L_{j}=\{i\mid\text{$e_{j}$ is an edge of~{}$\mathcal{H}_{i}$}\}.$
For each $1\leq i\leq k$ and each index $j$ such that $i\in L_{j}$, we choose an element $v_{i,j}\in e_{j}\setminus\cup_{\ell<j}e_{\ell}$. Recall that $i\in L_{j}$ means that $e_{j}$ is a circuit in $\mathcal{M}_{i}$. Hence, by  Observation 8.18, we can delete $(v_{i,j})_{\{j:i\in L_{j}\}}$ one by one, in descending order of $j$, to obtain:
$span_{\mathcal{M}_{i}}(\cup_{j:i\in L_{j}}e_{j}\cup\cup_{j:i\not\in L_{j}}e_{j})=span_{\mathcal{M}_{i}}\Big{(}\cup_{j:i\in L_{j}}(e_{j}\setminus\{v_{i,j}\})\cup\cup_{j:i\not\in L_{j}}e_{j}\Big{)}$
and then
$span_{\mathcal{M}_{i}}(\bigcup\mathcal{K})$
$=$ $span_{\mathcal{M}_{i}}(\cup_{j:i\in L_{j}}e_{j}\cup\cup_{j:i\not\in L_{j}}e_{j})$
$=$ $span_{\mathcal{M}_{i}}\Big{(}\cup_{j:i\in L_{j}}(e_{j}\setminus\{v_{i,j}\})\cup\cup_{j:i\not\in L_{j}}e_{j}\Big{)}$
$=$ $span_{\mathcal{M}_{i}}\Big{(}\cup_{j:i\in L_{j}}\big{(}e_{j}\setminus(\{v_{i,j}\}\cup\cup_{\ell<j}e_{\ell})\big{)}\cup\cup_{j:i\not\in L_{j}}(e_{j}\setminus\cup_{\ell<j}e_{\ell})\Big{)},$
which has size at most
(19) $t\Big{(}\sum_{j:i\in L_{j}}(|e_{j}\setminus\cup_{\ell<j}e_{\ell}|-1)+\sum_{j:i\not\in L_{j}}|e_{j}\setminus\cup_{\ell<j}e_{\ell}|\Big{)}.$
Let $f_{j}=e_{j}\setminus\cup_{\ell<j}e_{\ell}$. By the frugality, $|f_{j}|=|e_{j}\setminus\cup_{\ell<j}e_{\ell}|\geq 2$, so that
(20) $|f_{j}|\leq 2(|f_{j}|-1).$
Since for each $j$, $e_{j}\in\mathcal{H}_{i}$ for some $i\in L_{j}$, there are at most $k-1$ many indices $i^{\prime}$ in $[k]$ such that $i^{\prime}\not\in L_{j}$ Then by (18), (19), and (20) we have
$|V|$ $\leq t\sum_{i=1}^{k}\Big{(}\sum_{j:i\in L_{j}}(|f_{j}|-1)+\sum_{j:i\not\in L_{j}}|f_{j}|\Big{)}$
$\leq t\sum_{i=1}^{k}\Big{(}\sum_{j:i\in L_{j}}(|f_{j}|-1)+\sum_{j:i\not\in L_{j}}2(|f_{j}|-1)\Big{)}$
$=t\sum_{j=1}^{p}\Big{(}\sum_{i:i\in L_{j}}(|f_{j}|-1)+\sum_{i:i\not\in L_{j}}2(|f_{j}|-1)\Big{)}$
$\leq t\sum_{j=1}^{p}\Big{(}(|f_{j}|-1)+(k-1)2(|f_{j}|-1)\Big{)}$
$=(2k-1)t\sum_{j=1}^{p}(|f_{j}|-1).$
On the other hand, by (17)
$\eta(\mathcal{C})=\eta(\mathcal{I}(\mathcal{H}))\geq\gamma^{E}(\mathcal{H})=\sum_{j=1}^{p}(|e_{j}\setminus\cup_{\ell<j}e_{\ell}|-1)=\sum_{j=1}^{p}(|f_{j}|-1).$
We have thus shown
$\frac{|V(\mathcal{C})|}{\eta(\mathcal{C})}\leq(2k-1)t.$
The above argument works for any $S\subseteq V$ by considering the induced subhypergraph $\mathcal{H}[S]$, hence
$\Delta_{\eta}(\mathcal{C})=\max_{\emptyset\neq S\subseteq V}\frac{|S|}{\eta(\mathcal{C}[S])}\leq(2k-1)t,$
which completes the proof of the theorem. ∎
9. Fractional colorings and polytopes
Like other combinatorial concepts, “coloring” has a fractional version.
Definition 9.1.
Given a complex $\mathcal{C}$ on $V$, a fractional coloring of $\mathcal{C}$ is a function $f:\mathcal{C}\rightarrow\mathbb{R}_{\geq 0}$ satisfying $\sum_{S\in\mathcal{C}:v\in S}f(S)\geq 1$ for every $v\in V$. The fractional chromatic number $\chi^{*}(\mathcal{C})$ of $\mathcal{C}$ is minimum of $\sum_{S\in\mathcal{C}}f(S)$ over all fractional colorings $f$ of $\mathcal{C}$.
We shall also need a generalized weighted version:
Definition 9.2.
Given a complex $\mathcal{C}$ on $V$ and $\vec{h}\in\mathbb{R}^{V}_{\geq 0}$, an $\vec{h}$-fractional coloring of $\mathcal{C}$ is a function $f:\mathcal{C}\rightarrow\mathbb{R}_{\geq 0}$ satisfying $\sum_{S\in\mathcal{C}:v\in S}f(S)\geq h(v)$ for every $v\in V$. The fractional $\vec{h}$-chromatic number, denoted by $\chi^{*}(\mathcal{C},\vec{h})$, is the minimum of $\sum_{S\in\mathcal{C}}f(S)$ over all $\vec{h}$-fractional colorings $f$ of $\mathcal{C}$.
So, $\chi^{*}(\mathcal{C})=\chi^{*}(\mathcal{C},\vec{1})$.
The fractional $\vec{h}$-chromatic number is closely related to the $P$-polytope of $\mathcal{C}$, defined as:
Definition 9.3.
For a complex $\mathcal{C}$, let $P(\mathcal{C})$ be the convex hull $conv(\{\mathbf{1}_{S}\in\mathbb{R}^{V(\mathcal{C})}_{\geq 0}\mid S\in\mathcal{C}\})$.
Definition 9.4.
A polytope $Z\subseteq\mathbb{R}^{V}_{\geq 0}$ is called closed-down if $z\in Z$ implies that $y\in Z$ for every $0\leq y\leq z$.
Clearly, $P(\mathcal{C})$ is closed-down.
The two notions, $P(\mathcal{C})$ and the fractional $\vec{h}$-chromatic number, are linked by the following function:
Definition 9.5.
For a closed-down polytope $Z\subseteq\mathbb{R}^{V}_{\geq 0}$ and a vector $\vec{h}\in\mathbb{R}^{V}_{\geq 0}$, let
$\psi(Z,\vec{h}):=\min(\{t\in\mathbb{R}^{+}\mid\vec{h}/t\in Z\}).$
Theorem 9.6.
For any complex $\mathcal{C}$ and vector $\vec{h}\in\mathbb{R}_{\geq 0}^{V(\mathcal{C})}$,
$\chi^{*}(\mathcal{C},\vec{h})=\psi(P(\mathcal{C}),\vec{h}).$
Proof.
Let $\psi(P(\mathcal{C}),\vec{h})=t$. Then $\frac{\vec{h}}{t}\in P(\mathcal{C})$ implies that there exists $\lambda_{S}\geq 0$ for all $S\in\mathcal{C}$ so that $\vec{h}=\sum_{S\in\mathcal{C}}t\lambda_{S}\mathbf{1}_{S}$, where $\sum_{S\in\mathcal{C}}\lambda_{S}=1$. Then, the function $f:\mathcal{C}\rightarrow\mathbb{R}_{\geq 0}$ defined as $f(S):=t\lambda_{S}\geq 0$ satisfies that for each $v\in V(\mathcal{C})$, $\sum_{S\in\mathcal{C}:v\in S}f(S)=\sum_{S\in\mathcal{C}}t\lambda_{S}\mathbf{1}_{S}(v)=h(v)$. Thus $f$ is an $\vec{h}$-fractional coloring. Since $\sum_{S\in\mathcal{C}}f(S)=\sum_{S\in\mathcal{C}}t\lambda_{S}=t$, we have $\chi^{*}(\mathcal{C},\vec{h})\leq t=\psi(P(\mathcal{C}),\vec{h})$.
For the other direction, let $f:\mathcal{C}\rightarrow\mathbb{R}_{\geq 0}$ be an $\vec{h}$-fractional coloring satisfying $\sum_{S\in\mathcal{C}}f(S)=\chi^{*}(\mathcal{C},\vec{h})=a$. Let $\lambda_{S}=\frac{1}{a}f(S)$ for every $S\in\mathcal{C}$, and consider the function $g=\sum_{S\in\mathcal{C}}\lambda_{S}\mathbf{1}_{S}:V(\mathcal{C})\rightarrow\mathbb{R}_{\geq 0}$. This is a convex combination of the $(\mathbf{1}_{S})_{S\in\mathcal{C}}$ and thus $\vec{g}\in P(\mathcal{C})$. Since $f$ is an $\vec{h}$-fractional coloring, $g(v)=\sum_{S\in\mathcal{C}:v\in S}\lambda_{S}=\sum_{S\in\mathcal{C}:v\in S}\frac{1}{a}f(S)\geq\frac{h(v)}{a}$ for each $v\in V(\mathcal{C})$. So, $0\leq\frac{\vec{h}}{a}\leq\vec{g}$. Since $P(\mathcal{C})$ is closed-down, we have $\frac{\vec{h}}{a}\in P(\mathcal{C})$ and thus, $\psi(P(\mathcal{C}),\vec{h})\leq a=\chi^{*}(\mathcal{C},\vec{h})$. ∎
The theorem says that $\chi^{*}(\mathcal{C},\vec{h})$ is the first time $t$ you enter $P(\mathcal{C})$ along the direction $\vec{h}/t$, going towards the origin. On the other hand, for =$\mathcal{L}=\{\mathcal{M}_{1},\dots,\mathcal{M}_{k}\}\in\mathcal{D}^{k}$, to enter $\cap_{i=1}^{k}P(\mathcal{M}_{i})$ along the direction $\vec{h}$ towards the origin, you need to enter each  $P(\mathcal{M}_{i})$. The next is a rigorous formulation based on this fact.
Corollary 9.7.
Let =$\mathcal{L}=\{\mathcal{M}_{1},\cdots,\mathcal{M}_{k}\}\in\mathcal{D}^{k}$ on $V$ and let $\mathcal{C}=\bigcap\mathcal{L}$. Let $\lambda$ be a positive real number. The following are equivalent.
  1. (1)
    $\lambda P(\mathcal{C})\supseteq\cap_{i=1}^{k}P(\mathcal{M}_{i})$.
  2. (2)
    $\chi^{*}(\mathcal{C},\vec{h})\leq\lambda\max_{1\leq i\leq k}\chi^{*}(\mathcal{M}_{i},\vec{h})$ for every $h\in\mathbb{R}^{V}_{\geq 0}$.
Proof.
By Theorem 9.6, for any complex $\mathcal{C}$ and $\vec{h}\in\mathbb{R}^{V(C)}_{\geq 0}$, $\chi^{*}(\mathcal{C},\vec{h})=\psi(P(\mathcal{C}),\vec{h})$.
If (1) holds, let
$t=\max_{1\leq i\leq k}\chi^{*}(\mathcal{M}_{i},\vec{h})=\max_{1\leq i\leq k}\psi(P(\mathcal{M}_{i}),\vec{h}).$
Then by definition of $\psi$, $\frac{\vec{h}}{t}\in P(\mathcal{M}_{i})$ for each $i$, i.e., $\frac{\vec{h}}{t}\in\cap_{i=1}^{k}P(\mathcal{M}_{i})$. Then by the assumption, $\frac{\vec{h}}{\lambda t}\in P(\bigcap_{i=1}^{k}\mathcal{M}_{i})$ so that
$\chi^{*}(\cap_{i=1}^{k}\mathcal{M}_{i},\vec{h})=\psi\Big{(}P(\cap_{i=1}^{k}\mathcal{M}_{i}),\vec{h}\Big{)}\leq\lambda t.$
If (2) holds, for any $\vec{h}\in\cap_{i=1}^{k}P(\mathcal{M}_{i})$,
$\max_{1\leq i\leq k}\psi(P(\mathcal{M}_{i}),\vec{h})=\max_{1\leq i\leq k}\chi^{*}\Big{(}P(\mathcal{M}_{i}),\vec{h}\Big{)}\leq 1,$
which by the assumption implies
$\chi^{*}(\cap_{i=1}^{k}\mathcal{M}_{i},\vec{h})=\psi\Big{(}P(\cap_{i=1}^{k}\mathcal{M}_{i}),\vec{h}\Big{)}\leq\lambda.$
Therefore $\frac{\vec{h}}{\lambda}\in P(\cap_{i=1}^{k}\mathcal{M}_{i})$, which implies $\lambda P(\cap_{i=1}^{k}\mathcal{M}_{i})\supseteq\cap_{i=1}^{k}P(\mathcal{M}_{i})$. ∎
Recall Edmonds’ theorem:
Theorem 9.8.
For any pair of matroids $\mathcal{M}$ and $\mathcal{N}$ on the same ground set
(21) $P(\mathcal{M}\cap\mathcal{N})=P(\mathcal{M})\cap P(\mathcal{N})$
The equality (21) means that going in any direction $\vec{h}$, we enter $P(\mathcal{M}\cap\mathcal{N})$ precisely when we enter both of $P(\mathcal{M})$ and $P(\mathcal{N})$. Hence, by Corollary 9.7, Theorem 9.8 is equivalent to the following: for every $\vec{h}\in\mathbb{R}^{V}_{\geq 0}$,
(22) $\chi^{*}(\mathcal{M}\cap\mathcal{N},\vec{h})=\max(\chi^{*}(\mathcal{M},\vec{h}),\chi^{*}(\mathcal{N},\vec{h})).$
The case $\vec{h}=\vec{1}$ was studied in [2].
9.1. More on the fractional $\vec{h}$-chromatic number: $\chi^{*}$ vs. $\Delta$
We have seen that in a general complex $\mathcal{C}$ there holds $\chi(\mathcal{C})\leq\lceil\Delta(\mathcal{C})\rceil$. In [2] it was shown that in the case of the fractional chromatic number the ceilings can be dropped.
Theorem 9.9.
[2] For a complex $\mathcal{C}$, $\chi^{*}(\mathcal{C})\leq\Delta(\mathcal{C})$. When $\mathcal{C}$ is a matroid, equality holds.
This, too, can be generalized to weighted fractional versions.
Definition 9.10.
For a complex $\mathcal{C}$ and $\vec{h}\in\mathbb{R}^{V}_{\geq 0}$, let
$\Delta(\mathcal{C},\vec{h}):=\max_{\emptyset\neq S\subseteq V(\mathcal{C})}\frac{h[S]}{\min\Big{(}\eta(\mathcal{C}[S]),\;rank_{\mathcal{C}}(S)\Big{)}}.$
The afore-defined $\Delta(\mathcal{C})$ is the special case $\vec{h}=\vec{1}$.
Remark 9.11.
For a loopless matroid $\mathcal{M}$,
$\Delta(\mathcal{M},\vec{h})=\max_{\emptyset\neq S\subseteq V(\mathcal{M})}\frac{h[S]}{rank_{\mathcal{M}}(S)}.$
Theorem 9.9 is generalized as follows:
Theorem 9.12.
For any complex $\mathcal{C}$ and every $\vec{h}\in\mathbb{R}^{V}_{\geq 0}$, $\chi^{*}(\mathcal{C},\vec{h})\leq\Delta(\mathcal{C},\vec{h})$.
Proof.
Since any irrational number can be arbitrarily approximated by rational numbers and for any constant $t$, $\chi^{*}(\mathcal{C},t\vec{h})=t\chi^{*}(\mathcal{C},\vec{h})$ and $\Delta(\mathcal{C},t\vec{h})=t\Delta(\mathcal{C},\vec{h})$, we may assume that $\vec{h}$ is integral, thus $\Delta(\mathcal{M},\vec{h})=\frac{p}{q}$ for integers $p\geq 0$ and $q>0$. Furthermore we may assume that for every  $v\in V(\mathcal{C})$, $rank_{\mathcal{C}}(\{v\})=1$, since otherwise $\Delta(\mathcal{C},\vec{h})=\infty$ and the inequality holds.
Then let $W=\bigsqcup_{p}V(\mathcal{C})$ be the set of $p$ disjoint copies of $V(\mathcal{C})$. We consider the join complex $\mathcal{D}=*_{p}\mathcal{C}$ on $W$, and we consider a generalized partition matroid $\mathcal{N}$ on $W$ defined in the following way: for each $v\in V(\mathcal{C})$, the set $W_{v}$ consisting of the $p$ copies of $v$ in $W$ is a part, and the respective constraint parameter is $h(v)q$. Since $h(v)=\frac{h(v)}{rank_{\mathcal{C}}(\{v\})}\leq\Delta(\mathcal{C},\vec{h})=\frac{p}{q}$, we have $h(v)q\leq p$. Thus $\mathcal{N}$ is well-defined.
If there is a base $S$ of $\mathcal{N}$ that is in $\mathcal{D}$, then $S$ corresponds to $p$ faces $S_{1},\dots,S_{p}\in\mathcal{C}$ such that each $v\in V(\mathcal{C})$ is in $h(v)q$ of them. Then putting weight $\frac{1}{q}$ on each of the faces $S_{1},\dots,S_{p}$ gives an $\vec{h}$-fractional coloring of total weight $\frac{p}{q}=\Delta(\mathcal{C},\vec{h})$, which completes the proof.
It remains to prove there is a base of $\mathcal{N}$ that is in $\mathcal{D}$. For the matroid $\mathcal{N}$, the dual matroid $\mathcal{N}^{*}$ has the same ground set as $\mathcal{N}$, and a base of $\mathcal{N}$ is the complement of a base of $\mathcal{N}$. Therefore $\mathcal{N}^{*}$ is also a generalized partition matroid with parts $(W_{v})_{v\in V(\mathcal{C})}$ and respective constraint parameters $(p-h(v)q)_{v\in V(\mathcal{C})}$.
Let $U=\bigsqcup_{2}W$ be two disjoint copies of $W$, and we consider $\mathcal{D}*\mathcal{N}^{*}$ on $U$. For each $w\in W$, let $U_{w}$ be the set of two copies of $w$ in $U$. If we can show that there exists a choice function $\phi:W\rightarrow\cup_{w\in W}U_{w}$ such that $\phi(w)\in U_{w}$ for each $w\in W$ and $Im(\phi)\in\mathcal{D}*\mathcal{N}^{*}$, then it corresponds to a set $A\subseteq W$ such that $A\in\mathcal{D}$ and $W\setminus A\in\mathcal{N}^{*}$. It implies that $A$ contains a set $S$ that is in $\mathcal{D}$ and is a base in $\mathcal{N}$. We have done.
To show the existence of the choice function $\phi$, we shall apply Topological Hall (Theorem 4.11). And the condition to check is for every $X\subseteq W$,
$\eta\Big{(}(\mathcal{D}*\mathcal{N}^{*})[\cup_{w\in X}U_{w}]\Big{)}\geq|X|.$
By the super-additivity (Proposition 4.7) of $\eta$ with respect to the join operation,
$\eta\Big{(}(\mathcal{D}*\mathcal{N}^{*})[\cup_{w\in X}U_{w}]\Big{)}\geq\eta(\mathcal{D}[X])+\eta(\mathcal{N}^{*}[X])$
so that it is enough to show that for every $X\subseteq W$,
(23) $\eta(\mathcal{D}[X])+\eta(\mathcal{N}^{*}[X])\geq|X|.$
Fix some $X\subseteq W$. For each $v\in V(\mathcal{C})$, let
#$\#v=|X\cap W_{v}|$
be the number of copies of $v$ occurring in $X$. Clearly $0\leq\#v\leq p$ and $\sum_{v\in V(\mathcal{C})}\#v=|X|$. For each $1\leq i\leq p$, let $X_{i}$ be the intersection of $X$ with the $i$th copy of $V(\mathcal{C})$. Therefore $X$ is the disjoint union of $X_{i}$’s. By the definition of $\Delta(\mathcal{C},h)$, we have
$\frac{p}{q}=\Delta(\mathcal{C},h)\geq\frac{h[X_{i}]}{\eta(\mathcal{C}[X_{i}])}$
so that $\eta(\mathcal{C}[X_{i}])\geq\frac{q}{p}h[X_{i}]$. Again by the super-additivity (Proposition 4.7),
(24) $\eta(\mathcal{D}[X])\geq\sum_{i=1}^{p}\eta(\mathcal{C}[X_{i}])\geq\sum_{i=1}^{p}\frac{q}{p}h[X_{i}]=\sum_{v\in V(\mathcal{C})}\frac{q}{p}\#v\cdot h(v).$
On the other hand, by Proposition 4.8
(25) $\begin{split}\eta(\mathcal{N}^{*}[X])\geq rank_{\mathcal{N}^{*}}(X)&=\sum_{v\in V(\mathcal{C})}\min\Big{(}p-h(v)q,\;\#v\Big{)}\\ &\geq\sum_{v\in V(\mathcal{C})}(p-h(v)q)\frac{\#v}{p}.\end{split}$
Combining (24) and (25), we have
$\eta(\mathcal{D}[X])+\eta(\mathcal{N}^{*}[X])\geq\sum_{v\in V(\mathcal{C})}\#v=|X|,$
which proves (23). ∎
Theorem 9.13.
For any matroid $\mathcal{M}$ and $\vec{h}\in\mathbb{R}_{\geq 0}^{V(\mathcal{M})}$, $\chi^{*}(\mathcal{M},\vec{h})=\Delta(\mathcal{M},\vec{h})$.
Proof.
Combining with Theorem 9.12, it remains to show that $\chi^{*}(\mathcal{M},\vec{h})\geq\Delta(\mathcal{M},\vec{h})$ for loopless $\mathcal{M}$. Let $X\subseteq V$ attain $\frac{h[X]}{rank_{\mathcal{M}}(X)}=\Delta(\mathcal{M},\vec{h})$. Let $f$ be any $\vec{h}$-fractional coloring of $\mathcal{M}$. Then
$\sum_{v\in X}\sum_{S\in\mathcal{M}:v\in S}f(S)\geq h[X].$
On the other hand,
$\sum_{v\in X}\sum_{S\in\mathcal{M}:v\in S}f(S)\leq\sum_{S\in\mathcal{M}}\sum_{v\in X\cap S}f(S)\leq rank_{\mathcal{M}}(X)\sum_{S\in\mathcal{M}}f(S),$
which implies $\sum_{S\in\mathcal{M}}f(S)\geq\frac{h[X]}{rank_{\mathcal{M}}(S)}=\Delta(\mathcal{M},\vec{h})$. ∎
10. Two more polytopes associated with a complex
Definition 10.1.
  1. (1)
    Given a complex $\mathcal{C}$, let
    $Q(\mathcal{C})=\{f\in\mathbb{R}^{V(\mathcal{C})}_{\geq 0}\mid\sum_{v\in S}f(v)\leq rank_{\mathcal{C}}(S)\text{ for every }S\subseteq V(\mathcal{C})\}.$
  2. (2)
    For =$\mathcal{L}=\{\mathcal{M}_{1},\dots,\mathcal{M}_{k}\}\in\mathcal{D}^{k}$ let [^1] For a matroid $\mathcal{M}$, a theorem of Edmonds [12] is $P(\mathcal{M})=Q(\mathcal{M})$.
    $R(\mathcal{L})=\cap_{i=1}^{k}Q(\mathcal{M}_{i})=\cap_{i=1}^{k}P(\mathcal{M}_{i}).$
Clearly,
(26) $P(\mathcal{C})\subseteq Q(\mathcal{C})\subseteq R(\mathcal{L}).$
Theorem 9.8 says that for $k=2$ there is equality throughout. For $k>2$ both containments may be strict. (See Observation 10.2, Example 10.3, and Example 14.10.)
Observation 10.2.
For any constant $\lambda>0$, there exists a complex $\mathcal{C}$ such that $\lambda P(\mathcal{C})\not\supseteq Q(\mathcal{C})$.
Proof.
We construct a complex $\mathcal{C}$ and a vector $v\in\mathbb{R}^{V}_{\geq 0}$ such that $v\not\in\lambda P(\mathcal{C})$ but $v\in Q(\mathcal{C})$.
Let $k\in\mathbb{Z}^{+}$ be such that $\frac{1}{2}+\frac{1}{3}+\cdots+\frac{1}{k}>\lambda$, and let $n=2+3+\cdots+k$. We construct the complex $\mathcal{C}$ on $V=[n]$ in the following way. Let
$v=(\frac{1}{2},\frac{1}{2},\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4},...,\frac{1}{k},...,\frac{1}{k})\in\mathbb{R}^{n}.$
Let
$\mathcal{C}=\{A\subseteq[n]:v\cdot\mathbf{1}_{A}\leq 1\}.$
Since any vector $u\in P(\mathcal{C})$ is a convex combination of $\{\mathbf{1}_{A}\mid A\in\mathcal{C}\}$, then $v\cdot u\leq 1$. Therefore $u\in\lambda P(\mathcal{C})$ implies $v\cdot u\leq\lambda$. Since
$v\cdot v=\sum_{i=2}^{k}\frac{1}{i}\cdot\frac{1}{i}\cdot i>\lambda,$
we have $v\not\in\lambda P(\mathcal{C})$.
On the other hand, for $S\subseteq[n]$, let $r$ the minimal integer such that $-1<v[S]\leq r$. Then $S$ must have at least $r$ members with $v$-value at most $\frac{1}{r}$: if not,
$v[S]=v[\{j\in S:v(j)>1/r\}]+v[\{j\in S:v(j)\leq 1/r\}]\leq\sum_{i=2}^{r-1}\frac{1}{i}\cdot i+r\frac{1}{r}\leq r-1,$
a contradiction. Therefore the $r$ members form a face of $\mathcal{C}$ and we have $rank_{\mathcal{C}}(S)\geq r\geq v[S]$, which implies $v\in Q(\mathcal{C})$. ∎
$P(\mathcal{C})\neq Q(\mathcal{C})$ is possible even in flag complexes, as the following example shows.
Example 10.3.
Define a complex $\mathcal{C}$ on $V=\{x_{1},x_{2},\dots,x_{9},y_{1},y_{2},y_{3},z_{1},z_{2},z_{3}\}$ and the maximal faces of $\mathcal{C}$ are $\{y_{1},y_{2},y_{3}\},\{z_{1},z_{2},z_{3}\}$, and $\{y_{i},x_{3(i-1)+j}\}$, $\{z_{j},x_{3(i-1)+j}\}$ for $1\leq i,j\leq 3$. It can be verified that this complex is 2-determined so it is a flag complex. Therefore by Observation 3.7, $\mathcal{C}$ is the intersection of partition matroids.
Consider the vector $w$ such that $w(x_{i})=\frac{1}{9}$ for each $1\leq i\leq 9$ and $w(y_{j})=w(z_{j})=\frac{1}{4}$ for each $1\leq j\leq 9$. Then we claim that
$w\in Q(\mathcal{C})\setminus P(\mathcal{C}).$
To see that $w\in Q(\mathcal{C})$, for every $U\subseteq V$ such that $rank_{\mathcal{C}}(U)=1$, the maximal such $U$ in containment relation must be $\{x_{t}\}_{1\leq t\leq 9}$ or consist of four elements of $\{x_{t}\}_{1\leq t\leq 9}$, one elements of $\{y_{1},y_{2},y_{3}\}$, and one element of $\{z_{1},z_{2},z_{3}\}$. In either case we have $w[U]\leq 1$. Since $rank(\mathcal{C})=3$, for every set $U$ with $rank_{\mathcal{C}}(U)=3$, we have $w[U]\leq w[V]\leq\frac{1}{9}\cdot 9+\frac{1}{4}\cdot 6\leq 3$. The maximal rank 2 set in $\mathcal{C}$ has the form
$\{y_{i},y_{i^{\prime}},z_{j},z_{j^{\prime}},x_{1},x_{2},\dots,x_{9}\}$
for $i\neq i^{\prime}$ and $j\neq j^{\prime}$, which has $w$-weight $\frac{1}{9}\cdot 9+\frac{1}{4}\cdot 4\leq 2$, which implies for any $U$ with $rank_{\mathcal{C}}(U)=2$, $w[U]\leq 2$. Therefore $w\in Q(\mathcal{C})$.
To see that $w\not\in P(\mathcal{C})$, we are going to prove that $w$ cannot be written as a convex combination $\sum_{\lambda_{i}}\mathbf{1}_{S_{i}}$ for $S_{i}\in\mathcal{C}$. First, the coefficients of $\{y_{1},y_{2},y_{3}\}$ and $\{z_{1},z_{2},z_{3}\}$ cannot be positive, since the coefficients of the faces containing any $x_{i}$ for $1\leq i\leq 9$ should sum up to $\frac{1}{9}\cdot 9=1$. Second, for the sum of the coefficients of the faces of the form $\{x_{p},y_{i}\}$ or $\{x_{q},z_{j}\}$, when considering from the $\{x_{t}\}_{1\leq t\leq 9}$ side, the sum is at most $\frac{1}{9}\cdot 9=1$, but it should be at least $\frac{1}{4}\cdot 6=\frac{3}{2}$ when considering from the $\{y_{1},y_{2},y_{3},z_{1},z_{2},z_{3}\}$ side, which is impossible. Therefore $w\not\in P(\mathcal{C})$.
11. Bounding the ratios between the polytopes
Definition 11.1.
For two closed-down polytopes $A,B$ in $\mathbb{R}^{V}_{\geq 0}$ let
$B:A=\min\{t\in\mathbb{R}_{\geq 0}\mid tA\supseteq B\}$
(here, as usual, $\min\emptyset=\infty$).
We have seen that for $k>2$ the three complexes, $P(\mathcal{C}),Q(\mathcal{C})$ and $R(\mathcal{L})$ (where $\mathcal{L}\in\mathcal{D}^{k}$ and $\mathcal{C}=\bigcap\mathcal{L}$), may differ. But to what extent? That is, what are the maximal values of $R(\mathcal{L}):Q(\mathcal{C})$, $R(\mathcal{L}):P(\mathcal{C})$ and $Q(\mathcal{C}):P(\mathcal{C})$? These are matroidal counterparts of questions studied on fractional matchings in hypergraphs, we aim to connect the two types of questions.
Theorem 11.2.
For any $\mathcal{L}\in\mathcal{D}^{k}$,
$R(\mathcal{L}):P(\bigcap\mathcal{L})\leq k.$
By Corollary 9.7 the theorem is equivalent to:
Theorem 11.3.
Let =$\mathcal{L}=\{\mathcal{M}_{1},\cdots,\mathcal{M}_{k}\}\in\mathcal{D}^{k}$, $\mathcal{C}=\bigcap\mathcal{L}$, and $\vec{h}\in\mathbb{R}^{V}_{\geq 0}$. Then
$\chi^{*}(\mathcal{C},\vec{h})\leq k\max_{1\leq i\leq k}\chi^{*}(\mathcal{M}_{i},\vec{h}).$
The tools needed for the proof are developed in the next sections.
12. Weighted matroids and polytopes
Arguably, the most basic combinatorial min-max result is that in a matroid the rank (maximal size of an independent set) equals the minimal size of a spanning set. This has a weighted version, in which every vertex $v$ has a weight $w(v)$. The definition of a maximum weight independent set is natural — the maximal sum of weights in an independent set. This is denoted by $\nu_{w}$. But what about weighted spanning sets? A crucial observation is that a “weighted cover” does not span vertices, but their weights. While the matroid is defined on sets of vertices, weighted spanning is defined on functions. A dominating function is one that “spans” the given weight function $w$. This is captured by the following definition:
Notation 12.1.
For a matroid $\mathcal{M}$ and a function $f:V(\mathcal{M})\rightarrow\mathbb{R}$, for every $v\in V$, let
$f^{\mathcal{M}}(v)=\max_{T\subseteq V:v\in span_{\mathcal{M}}(T)}\min_{u\in T}f(u).$
Definition 12.2.
Given a matroid $\mathcal{M}$ and a weight function $w$ on $V(\mathcal{M})$, a function $f$ is called $w$-spanning if $f^{\mathcal{M}}\geq w$.
$f^{\mathcal{M}}$ is the function “spanned” by $f$. In particular, if $f=\mathbf{1}_{A}$ for some $A\subseteq V(\mathcal{M})$ then $f^{\mathcal{M}}=\mathbf{1}_{span_{\mathcal{M}}(A)}$.
Definition 12.3.
Let =$\mathcal{L}=\{\mathcal{M}_{1},\dots,\mathcal{M}_{k}\}\in\mathcal{D}^{k}$ on $V$. Given a weight function $\vec{w}\in\mathbb{R}^{V}_{\geq 0}$, a $k$-tuple $(f_{1},f_{2},\dots f_{k})$ of functions in $\mathbb{R}^{V}_{\geq 0}$ is a $\vec{w}$-fractional cover of $\mathcal{L}$ if $\sum_{i=1}^{k}f_{i}^{\mathcal{M}_{i}}(v)\geq w(v)$ for every $v\in V$. The $\vec{w}$-fractional covering number $\tau^{*}_{w}(\mathcal{L})$ is the minimum of $\sum_{i=1}^{k}f_{i}[V]$ over all $\vec{w}$-fractional covers $(f_{1},f_{2},\dots f_{k})$. The $\vec{w}$-covering number $\tau_{w}(\mathcal{L})$ is the minimum of $\sum_{i=1}^{k}f_{i}[V]$ over all $\vec{w}$-fractional covers $(f_{1},f_{2},\dots f_{k})$, where each $f_{i}$ has integral values.
Notation 12.4.
When $\vec{w}=\vec{1}$, these are called respectively fractional cover, fractional covering number, and covering number, and the subscript $\vec{1}$ is omitted.
Remark 12.5.
This deviates from the notation in [2], where our $\tau_{w}^{*}$ is denoted by $\tau_{w}$.
When the functions $f_{i}$ in a fractional cover $(f_{1},\dots,f_{k})$ are integral and $\sum_{i=1}^{k}f_{i}[V]$ is minimum, the functions $f_{i}$ must be 0-1 functions, and then so are $f_{i}^{\mathcal{M}_{i}}$, and $S_{i}:=f_{i}^{-1}(1)\in\mathcal{M}_{i}$. Furthermore, for $V_{i}:=(f^{\mathcal{M}_{i}})^{-1}(1)$, $f_{i}[V]=|S_{i}|=rank_{\mathcal{M}_{i}}(V_{i})$ and $\cup_{i=1}^{k}V_{i}=V$. Thus
$\tau(\mathcal{L})=\min_{V_{1}\cup\cdots\cup V_{k}=V}\sum_{i=1}^{k}rank_{\mathcal{M}_{i}}(V_{i}).$
Definition 12.6.
Let =$\mathcal{L}=\{\mathcal{M}_{1},\dots,\mathcal{M}_{k}\}\in\mathcal{D}^{k}$ on $V$. A vector $\vec{x}\in R(\mathcal{L})$ is called a matroidal fractional matching of $\mathcal{L}$. Given $\vec{w}\in\mathbb{R}^{V}_{\geq 0}$, the matroidal $\vec{w}$-fractional matching number $\nu^{*}_{w}(\mathcal{L})$ is the maximum of $\vec{x}\cdot\vec{w}$ over all matroidal fractional matchings $\vec{x}$. The $\vec{w}$-matching number $\nu_{w}(\mathcal{L})$ is the maximum of $\vec{x}\cdot\vec{w}$ over all matroidal fractional matchings $\vec{x}$ with integral coordinates.
Notation 12.7.
When $\vec{w}=\vec{1}$, these parameters are plainly called the matroidal fractional matching number and the matroidal matching number, respectively, and they are denoted by $\nu^{*}(\mathcal{L})$ and $\nu(\mathcal{L})$, respectively.
If a matroidal fractional matching $\vec{x}$ is integral, then $supp(\vec{x})\in\mathcal{M}_{1}\cap\cdots\cap\mathcal{M}_{k}$. Thus $\nu_{w}(\mathcal{L})$ is the maximum weight of a common independent set in $\mathcal{M}_{1}\cap\cdots\cap\mathcal{M}_{k}$. In particular,
$\nu(\mathcal{L})=rank(\mathcal{M}_{1}\cap\cdots\cap\mathcal{M}_{k}).$
The following theorem states that the two matroidal parameters $\tau^{*}_{w}$ and $\nu^{*}_{w}$ are equal.
Theorem 12.8.
For any =$\mathcal{L}=\{\mathcal{M}_{1},\dots,\mathcal{M}_{k}\}\in\mathcal{D}^{k}$ on $V$ and $w\in\mathbb{R}^{V}_{\geq 0}$,
$\nu^{*}_{w}(\mathcal{L})=\tau^{*}_{w}(\mathcal{L}).$
Proof.
Since a matroidal fractional matching $\vec{x}\in R(\mathcal{L})=\cap_{i=1}^{k}Q(\mathcal{M}_{i})$ satisfies $x\geq 0$ and $x[U]\leq rank_{\mathcal{M}_{i}}(U)$ for every $1\leq i\leq k$ and $U\subseteq V$, the dual LP of the matroidal $\vec{w}$-fractional matching number is
$\min\sum_{1\leq i\leq k,U\subseteq V}rank_{\mathcal{M}_{i}}(U)y_{i}(U)$
subject to
$\sum_{1\leq i\leq k,U\subseteq V:v\in U}y_{i}(U)\geq w(v)\text{ for every $v\in V$},$
$y_{i}(U)\geq 0\text{ for every $U\subseteq V$ and $1\leq i\leq k$.}$
The optimum value of this dual LP is equal to $\nu^{*}(\mathcal{L})$. It remains to show that the optimum value of this dual LP is equal to $\tau^{*}_{w}$.
First, for a $w$-fractional cover $(f_{1},\dots,f_{k})$ attaining $\tau_{w}^{*}$, assume $Im(f_{i})=\{a_{i,\ell}\}_{\ell}$ for $a_{i,1}>\cdots>a_{i,\ell_{i}}\geq 0$. Let $U_{i,\ell}=f^{-1}(a_{i,\ell})$. Then by the minimum property of $\sum_{i}f_{i}[V]$, $U_{i,\ell}\in\mathcal{M}_{i}$ for each $i$ and $1\leq\ell<\ell_{i}$. For each $i$, let $y_{i}(U)=a_{i,\ell}$ if $U=span_{\mathcal{M}_{i}}(U_{i,\ell})$ for some $1\leq\ell\leq\ell_{i}$ and $y_{i}(U)=0$ otherwise. Then $(y_{1},\dots,y_{k})$ defined in this way satisfies the constraints of the dual LP, and
$\tau^{*}_{w}=\sum_{i}f_{i}[V]=\sum_{i}\sum_{U}rank_{\mathcal{M}_{i}}(U)y_{i}(U).$
Therefore $\tau^{*}\geq\nu^{*}$.
On the other hand, for any optimum solution $(y_{1},\dots,y_{k})$ of the dual LP, let $\mathcal{S}_{i}\subseteq 2^{V}$ be the support of $y_{i}$ for $1\leq i\leq k$. For each $U\in\mathcal{S}_{i}$, take an $I(U)\in\mathcal{M}_{i}[U]$ satisfying $|I(U)|=rank_{\mathcal{M}_{i}}(U)$. Let $\mathcal{S}^{*}_{i}:=\{I(U)\mid U\in\mathcal{S}_{i}\}$. Set $y^{*}_{i}(Z)=y_{i}(U)$ if $Z=I(U)\in\mathcal{S}^{*}_{i}$ for some $U\in\mathcal{S}_{i}$ and $y^{*}_{i}(Z)=0$ if $Z\not\in\mathcal{S}_{i}^{*}$. Then
(27) $\sum_{1\leq i\leq k,U\subseteq V}rank_{\mathcal{M}_{i}}(U)y_{i}(U)=\sum_{1\leq i\leq k,U\subseteq V}rank_{\mathcal{M}_{i}}(U)y^{*}_{i}(U).$
Furthermore, we may assume for distinct $S,S^{\prime}\in\mathcal{S}^{*}_{i}$, $S\cap S^{\prime}=\emptyset$, otherwise we may define $z^{*}_{i}(S\setminus S^{\prime})=y^{*}_{i}(S)$, $z^{*}_{i}(S^{\prime}\setminus S)=y^{*}_{i}(S^{\prime})$, $z^{*}_{i}(S\cap S^{\prime})=y^{*}_{i}(S)+y^{*}_{i}(S^{\prime})$, and $z^{*}_{i}(T)=y^{*}_{i}(T)$ for all the other $T\subseteq V$. In this way, since $rank_{\mathcal{M}_{i}}(S)=rank_{\mathcal{M}_{i}}(S\setminus S^{\prime})+rank_{\mathcal{M}_{i}}(S\cap S^{\prime})$ and $rank_{\mathcal{M}_{i}}(S^{\prime})=rank_{\mathcal{M}_{i}}(S^{\prime}\setminus S)+rank_{\mathcal{M}_{i}}(S\cap S^{\prime})$, we have
(28) $\sum_{1\leq i\leq k,U\subseteq V}rank_{\mathcal{M}_{i}}(U)z^{*}_{i}(U)=\sum_{1\leq i\leq k,U\subseteq V}rank_{\mathcal{M}_{i}}(U)y^{*}_{i}(U)$
and suppose $S=I(U)$, we still have $U\subseteq span_{\mathcal{M}_{i}}((S\setminus S^{\prime})\cup(S\cap S^{\prime}))$ and $y_{i}(U)=y_{i}^{*}(S)\leq\min(z^{*}_{i}(S\setminus S^{\prime}),z^{*}_{i}(S\cap S^{\prime}))$. This decreases the number of the intersecting pairs and we can replace $y_{i}^{*}$ by $z_{i}^{*}$, and do it repeatedly until the sets in the support become mutually disjoint.
Then for each $S\in\mathcal{S}_{i}^{*}$, we set $f_{i}(v)=y_{i}^{*}(S)$ for each $v\in S$. The constraint in the dual LP and the above argument guarantee that $\sum_{i=1}^{k}f_{i}^{\mathcal{M}_{i}}(v)\geq w(v)$ for each $v\in V$, and using (28) and (27), we have
$\sum_{i=1}^{k}f_{i}[V]=\sum_{1\leq i\leq k,U\subseteq V}rank_{\mathcal{M}_{i}}(U)y^{*}_{i}(U)=\sum_{1\leq i\leq k,U\subseteq V}rank_{\mathcal{M}_{i}}(U)y_{i}(U).$
Therefore $\tau^{*}\leq\nu^{*}$. ∎
Corollary 12.9.
For any $\mathcal{L}\in\mathcal{D}^{k}$ on $V$ and $w\in\mathbb{R}^{V}_{\geq 0}$,
$\nu_{w}(\mathcal{L})\leq\nu^{*}_{w}(\mathcal{L})=\tau^{*}_{w}(\mathcal{L})\leq\tau_{w}(\mathcal{L}).$
Efforts have been made to bound the gap between these parameters. In [2, Conjecture 7.1] it was conjectured that for $\mathcal{L}\in\mathcal{D}^{k}$,
$\tau(\mathcal{L})\leq(k-1)\nu(\mathcal{L}).$
The following is a generalization to the weighted case.
Conjecture 12.10.
For any $\mathcal{L}\in\mathcal{D}^{k}$ on $V$ and $\vec{w}\in\mathbb{R}^{V}_{\geq 0}$,
$\tau_{w}^{*}(\mathcal{L})\leq(k-1)\nu_{w}(\mathcal{L}).$
The $(k-1)$-factor is best possible. See Example 14.10.
By Definition 12.6, the conjecture actually asks for the integrality gap of the LP $\tau^{*}_{w}(\mathcal{L})=\nu^{*}_{w}(\mathcal{L})=\max\vec{w}\cdot\vec{x}$ for $\vec{x}\in R(\mathcal{L})$. When =$\mathcal{L}=\{\mathcal{M}_{1},\mathcal{M}_{2}\}\in\mathcal{D}^{2}$, Edmonds’ matroid intersection theorem $R(\mathcal{L})=P(\mathcal{M}_{1})\cap P(\mathcal{M}_{2})=P(\mathcal{M}_{1}\cap\mathcal{M}_{2})$ proves that the polytope is integral, therefore for every $w\in\mathbb{R}^{V}_{\geq 0}$,
(29) $\tau_{w}^{*}(\mathcal{L})=\nu_{w}^{*}(\mathcal{L})=\nu_{w}(\mathcal{L}).$
For general $k$ we have:
Theorem 12.11.
For any $\mathcal{L}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{k})\in\mathcal{D}^{k}$ on $V$ and $\vec{w}\in\mathbb{R}^{V}_{\geq 0}$,
$\tau_{w}^{*}(\mathcal{L})\leq k\nu_{w}(\mathcal{L}).$
Proof.
Let $I\in\mathcal{M}_{1}\cap\cdots\cap\mathcal{M}_{k}$ satisfying $w[I]=\nu_{w}(\mathcal{L})$ and $w(u)>0$ for each $u\in I$. For each $1\leq i\leq k$, we set $f_{i}(x)=w(x)$ if $x\in I$ and $f_{i}(x)=0$ if $x\in V\setminus I$. Then $\sum_{i=1}^{k}f_{i}[V]=k\nu_{w}$. It remains to verify that $\sum_{i=1}^{k}f_{i}^{\mathcal{M}_{i}}(v)\geq w(v)$ for each $v\in V$. Suppose not. Then there exists $v\in V$ such that $\sum_{i=1}^{k}f_{i}^{\mathcal{M}_{i}}(v)<w(v)$. Since for $x\in I$, $\sum_{i=1}^{k}f_{i}^{\mathcal{M}_{i}}(x)\geq\sum_{i=1}^{k}f_{i}(x)=kw(x)$, this $v$ is not in $I$. Let $J$ be the set of indices $j$ such that $v\in span_{\mathcal{M}_{j}}(I)$. Then $\sum_{i=1}^{k}f_{i}^{\mathcal{M}_{i}}(v)=\sum_{j\in J}f_{j}^{\mathcal{M}_{j}}(v)<w(v)$. It means for each $j\in J$ there exists circuit $C_{j}$ in $\mathcal{M}_{j}$ and $u_{j}\in C_{j}\setminus\{v\}$ such that $v\in C_{j}$ and $C_{j}\setminus\{v\}\subseteq I$, which satisfies
(30) $\sum_{j\in J}w(u_{j})<w(v).$
[…]
$I^{\prime}:=I\cup\{v\}\setminus\{u_{j}\mid j\in J\}\in\cap_{i=1}^{k}\mathcal{M}_{i}=(\cap_{j\in J}\mathcal{M}_{j})\cap(\cap_{t\in[k]\setminus J}\mathcal{M}_{t}).$
But by (30), $w(I^{\prime})\geq w(I)-\sum_{j\in J}w(u_{j})+w(v)>w(I)$, contradictory to the maximum weight assumption of $I$. ∎
13. The ratio $R:Q$
Notation 13.1.
For =$\mathcal{L}=\{\mathcal{M}_{1},\dots,\mathcal{M}_{k}\}\in\mathcal{D}^{k}$ on $V$, for each $U\subseteq V$, let $\mathcal{L}_{U}$ be $\{\mathcal{N}_{1},\dots,\mathcal{N}_{k}\}$, where for each $1\leq i\leq k$$\mathcal{N}_{i}$ is the matroid with ground set $V$ and with the collection of independent set $\mathcal{M}_{i}[U]$.
Theorem 13.2.
For every $k$ and $\mathcal{L}\in\mathcal{D}^{k}$ on $V$,
$R(\mathcal{L}):Q(\bigcap\mathcal{L})=\max_{U:U\subseteq V}\nu^{*}(\mathcal{L}_{U}):\nu(\mathcal{L}_{U}).$
Proof.
Assume =$\mathcal{L}=\{\mathcal{M}_{1},\dots,\mathcal{M}_{k}\}$. Suppose $\lambda=\max_{U:U\subseteq V}\nu^{*}(\mathcal{L}_{U}):\nu(\mathcal{L}_{U})$. First we shall show that $R(\mathcal{L}):Q(\bigcap\mathcal{L})\leq\lambda$, which is equivalent to show that for every $x\in R(\mathcal{L})$, $x\in\lambda Q(\mathcal{C})$ for $\mathcal{C}=\bigcap\mathcal{L}$, i.e.,
(31) $x[U]\leq\lambda\cdot rank_{\mathcal{C}}(U)$
for every $U\subseteq V$.
Fix some $U\subseteq V$ and assume $\mathcal{L}_{U}=\{\mathcal{N}_{1},\dots,\mathcal{N}_{k}\}$, where the matroids $\mathcal{N}_{i}$s are as constructed in Notation 13.1. We define a vector $\bar{x}\in\mathbb{R}^{V}_{\geq 0}$ as the following: $\bar{x}(v)=x(v)$ for $v\in U$ and $\bar{x}(v)=0$ for $v\in V\setminus U$. Then $x\in R(\mathcal{L})=\cap_{i=1}^{k}Q(\mathcal{M}_{i})$ implies that for every $W\subseteq V$,
$\bar{x}[W]=x[W\cap U]\leq\min_{1\leq i\leq k}rank_{\mathcal{M}_{i}}(W\cap U)=\min_{1\leq i\leq k}rank_{\mathcal{N}_{i}}(W).$
Thus $\bar{x}\in R(\mathcal{L}_{U})=\cap_{i=1}^{k}Q(\mathcal{N}_{i})$. By the assumption of $\lambda$, $\nu^{*}(\mathcal{L}_{U})\leq\lambda\nu(\mathcal{L}_{U})=\lambda\cdot rank_{\mathcal{C}}(U)$, and then
$x[U]=\bar{x}[U]\leq\nu^{*}(\mathcal{L}_{U})\leq\lambda\nu(\mathcal{L}_{U})=\lambda\cdot rank_{\mathcal{C}}(U),$
as desired in (31).
On the other hand, suppose $R(\mathcal{L}):Q(\bigcap\mathcal{L})=\gamma$. We are going to show that $\max_{U:U\subseteq V}\nu^{*}(\mathcal{L}_{U}):\nu(\mathcal{L}_{U})\leq\gamma$, i.e.,
(32) $\nu^{*}(\mathcal{L}_{U})\leq\gamma\nu(\mathcal{L}_{U})$
for every $U\subseteq V$. Fix $U\subseteq V$ and let $x\in R(\mathcal{L}_{U})$ satisfying $x[V]=\nu^{*}(\mathcal{L}_{U})$. By the construction of $\mathcal{L}_{U}$, for every $W\subseteq V$,
$x[W]=x[W\cap U]\leq\min_{1\leq i\leq k}rank_{\mathcal{M}_{i}[U]}(W\cap U)\leq\min_{1\leq i\leq k}rank_{\mathcal{M}_{i}}(W),$
which implies $x\in R(\mathcal{L})=\cap_{i=1}^{k}Q(\mathcal{M}_{i})$. Then by the assumption, $\frac{x}{\gamma}\in Q(\bigcap\mathcal{L})$. Therefore for every $W\subseteq V$,
$\frac{x[W]}{\gamma}=\frac{x[W\cap U]}{\gamma}\leq rank_{\bigcap\mathcal{L}}(W\cap U)=rank_{\bigcap\mathcal{L}_{U}}(W),$
which implies $\frac{x}{\gamma}\in Q(\bigcap\mathcal{L}_{U})$. Therefore
$\frac{\nu^{*}(\mathcal{L}_{U})}{\gamma}=\frac{x[V]}{\gamma}\leq rank_{\cap\mathcal{L}_{U}}(V)=\nu(\mathcal{L}_{U}),$
as desired in (32). ∎
By Theorem 12.11 for any $\mathcal{L}\in\mathcal{D}^{k}$, $\nu^{*}(\mathcal{L})\leq k\nu(\mathcal{L})$, and hence we prove the following result.
Corollary 13.3.
For every $k$ and $\mathcal{L}\in\mathcal{D}^{k}$,
$R(\mathcal{L}):Q(\bigcap\mathcal{L})\leq k.$
This may well be superseded:
Conjecture 13.4.
For every $k$ and $\mathcal{L}\in\mathcal{D}^{k}$,
$R(\mathcal{L}):Q(\bigcap\mathcal{L})\leq k-1.$
We know this conjecture for $k=3$.
Corollary 13.5.
If $\mathcal{L}\in\mathcal{D}^{3}$ then $R(\mathcal{L}):Q(\bigcap\mathcal{L})\leq 2$.
Proof.
By Theorem 7.3 in [2] (a matroidal version of the 3-dimensional Ryser’s conjecture) $\nu(\mathcal{L}^{\prime})\geq\tau(\mathcal{L}^{\prime})/2$ for any $\mathcal{L}^{\prime}\in\mathcal{D}^{3}$. Since $\tau\geq\tau^{*}=\nu^{*}$ (Corollary 12.9), this together with Theorem 13.2 implies that $R(\mathcal{L}):Q(\bigcap\mathcal{L})\leq 2$. ∎
14. The ratio $R:P$
Mimicking the proof of [2, Lemma 9.2], one can prove that for any matroid $\mathcal{M}$ and $h,f\in\mathbb{R}_{\geq 0}^{V(\mathcal{M})}$,
$\sum_{v\in V(\mathcal{M})}h(v)f^{\mathcal{M}}(v)\leq\Delta(\mathcal{M},\vec{h})f[V].$
And mimicking the proof of [2, Theorem 10.4] yields that, for $\lambda>0$, if $\tau_{w}^{*}(\mathcal{L})\leq\lambda\nu_{w}(\mathcal{L})$ for every $w\in\mathbb{R}^{V}_{\geq 0}$, then $\lambda P(\bigcap\mathcal{L})\supseteq R(\mathcal{L})$, which in particular together with Theorem 12.11 implies $kP(\bigcap\mathcal{L})\supseteq R(\mathcal{L})$ as in Theorem 11.2. Here we give a more direct proof.
Theorem 14.1.
For every $k$ and $\mathcal{L}\in\mathcal{D}^{k}$ on $V$,
$R(\mathcal{L}):P(\bigcap\mathcal{L})=\max_{w\in\mathbb{R}^{V}_{\geq 0}}\tau_{w}^{*}(\mathcal{L}):\nu_{w}(\mathcal{L}).$
Proof.
Suppose $\lambda=\max_{w\in\mathbb{R}^{V}_{\geq 0}}\tau_{w}^{*}(\mathcal{L}):\nu_{w}(\mathcal{L})$. First we are going to show that
(33) $R(\mathcal{L}):P(\bigcap\mathcal{L})\leq\lambda.$
If not, i.e., $R(\mathcal{L})\setminus\lambda P(\bigcap\mathcal{L})\neq\emptyset$. Since $\lambda P$ is convex, it is the intersection of a finite set of halfspaces, i.e., there exist vectors $w_{1},\dots,w_{m}$ such that $x\in\lambda P$ if and only if for every $1\leq i\leq j$, $w_{i}\cdot x\leq a_{i}$ for some constant $a_{i}$ and for every $j<i\leq m$, $w_{i}\cdot x\geq b_{i}$. And since $\lambda P=\lambda\cdot conv(\{\mathbf{1}_{S}\mid S\in\bigcap\mathcal{L}\})$ is closed down and bounded, we may assume that $w_{i}\geq 0$ for every $1\leq i\leq m$. In particular, for $z\in R\setminus\lambda P$, since $R$ is closed down, the constraint violated must be the first type, i.e., $w_{i}\cdot z>a_{i}$ for some $1\leq i\leq j$. This implies $\tau^{*}_{w_{i}}(\mathcal{L})=\nu^{*}_{w_{i}}(\mathcal{L})>\lambda\nu_{w_{i}}(\mathcal{L})$, a contradiction to the choice of $\lambda$. Therefore (33) holds.
On the other hand, suppose $R(\mathcal{L}):P(\bigcap\mathcal{L})=\gamma$. We are going to show that
(34) $\max_{w\in\mathbb{R}^{V}_{\geq 0}}\tau_{w}^{*}(\mathcal{L}):\nu_{w}(\mathcal{L})\leq\gamma.$
For any $w\in\mathbb{R}^{V}_{\geq 0}$, let $x\in R(\mathcal{L})$ be a matroidal fractional matching satisfying $w\cdot x=\nu^{*}_{w}(\mathcal{L})$. Then by the assumption $\frac{x}{\gamma}\in P(\mathcal{C})$ for $\mathcal{C}=\bigcap\mathcal{L}$, which means there exists a collection of $\{S_{i}\in\mathcal{C}:i\in I\}$ such that $\frac{x}{\gamma}=\sum_{i\in I}\lambda_{i}\mathbf{1}_{S_{i}}$ is a convex combination. We take $S_{t}$ in the collection such that $\mathbf{1}_{S_{t}}\cdot w=\max_{i\in I}\mathbf{1}_{S_{i}}\cdot w$. Therefore
$\frac{\nu^{*}_{w}(\mathcal{L})}{\gamma}=\frac{x\cdot w}{\gamma}=\sum_{i\in I}\lambda_{i}\mathbf{1}_{S_{i}}\cdot w\leq\mathbf{1}_{S_{t}}\cdot w\leq\nu_{w}(\mathcal{L}),$
as desired in (34). ∎
Proof of Theorem 11.2.
The theorem follows from Theorem 12.11 and Theorem 14.1. ∎
14.1. The case of intersection of partition matroids
Over all we are doing hovers a well-known conjecture. The terms “$k-1$” in Conjecture 12.10 and its like are inspired by it. It relates to the intersection of partition matroids, but it can be formulated also for general matroids. It is Ryser’s conjecture (one of two famous “Ryser’s conjectures” — the other, regarding transversals in Latin squares, has recently been solved). To state the conjecture, we need the following definition.
Definition 14.2.
Let $\mathcal{H}$ be a hypergraph and $w:\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ be a function.
  1. (1)
    A fractional matching in the hypergraph $\mathcal{H}$ is a function $g:\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ satisfying
    $\sum_{e:v\in e}g(e)\leq 1$
    for every $v\in V(\mathcal{H})$. By $\nu^{*}_{w}(\mathcal{H})$ we denote the maximal value $\sum_{e\in\mathcal{H}}w(e)g(e)$ over all fractional matchings $g$. By $\nu_{w}(\mathcal{H})$ we denote the maximum of $\sum_{e\in\mathcal{H}}w(e)g(e)$ over all fractional matchings $g$ with integral values.
  2. (2)
    A fractional $w$-cover is a non-negative function $t$ on $V(\mathcal{H})$ satisfying
    $\sum_{v\in e}t(v)\geq w(e)$
    for all $e\in\mathcal{H}$. By $\tau^{*}_{w}(\mathcal{H})$ we denote the minimal $g[V]$ over all fraction $w$-covers $g$. By $\tau_{w}(\mathcal{H})$ we denote the minimum $g[V]$ over all fractional $w$-covers $g$ having integral values.
For $w\equiv 1$ we omit the subscript $w$ in both notations.
Remark 14.3.
We are using again the notation $\nu$ and $\tau$. It is related, but not identical, to our previous usage, in $\nu(\mathcal{L})$ and $\tau(\mathcal{L})$ defined in Section 12.
By LP duality $\tau^{*}_{w}(\mathcal{H})=\nu^{*}_{w}(\mathcal{H})$. We can now state Ryser’s conjecture, and we shall do it in the more general weighted version:
Conjecture 14.4 (Ryser’s conjecture).
Let $\mathcal{H}$ be a $k$-partite hypergraph, and let $w$ be a non-negative function on $E(\mathcal{H})$. Then $\tau_{w}(\mathcal{H})\leq(k-1)\nu_{w}(\mathcal{H})$.
The conjecture was proved in [1] in the case $k=3$ and $w\equiv 1$, using Theorem 4.11.
Füredi proved a fractional version for general $k$.
Theorem 14.5 (Füredi [14]).
In a $k$-partite hypergraph $\nu^{*}(\mathcal{H})\leq(k-1)\nu(\mathcal{H})$.
A topological proof is given in Appendix A. Füredi, Kahn and Seymour proved a weighted version of Füredi’s theorem:
Theorem 14.6 (Füredi–Kahn–Seymour [15]).
If $\mathcal{H}$ is $k$-partite then for every $w:\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$, $\nu^{*}_{w}(\mathcal{H})\leq(k-1)\nu_{w}(\mathcal{H})$.
Applying this result, we prove Conjecture 12.10 and thus also Conjecture 14.8 and Conjecture 14.9 below, for partition matroids.
Theorem 14.7.
If =$\mathcal{L}=\{\mathcal{M}_{1},\dots,\mathcal{M}_{k}\}$ on $V$ is $k$-tuple of partition matroids, then for every $w\in\mathbb{R}^{V}_{\geq 0}$,
$\tau_{w}^{*}(\mathcal{L})\leq(k-1)\nu_{w}(\mathcal{L}).$
Proof.
Let $(S_{i,j})_{1\leq j\leq p_{i}}$ be the parts of the partition matroids $\mathcal{M}_{i}$ for each $1\leq i\leq k$.
We construct a $k$-partite hypergraph $\mathcal{H}$ as the following: each part $S_{i,j}$ for $1\leq i\leq k$ and $1\leq j\leq p_{i}$ corresponds to a vertex of $\mathcal{H}$, which we still denote by $S_{i,j}$, and each $v\in\cap_{i=1}^{k}S_{i,\ell_{i}(v)}$ of the ground set of the matroids corresponds to an edge of $\mathcal{H}$, which is still denoted by $v$ and includes $S_{1,\ell_{1}(v)},\dots,S_{k,\ell_{k}(v)}$. Then such $\mathcal{H}$ is $k$-partite with parts $\cup_{1\leq j\leq p_{1}}S_{1,j},\dots,\cup_{1\leq j\leq p_{k}}S_{k,j}$.
Note that an $I\in\cap_{i=1}^{k}\mathcal{M}_{i}$ corresponds to a matching in $\mathcal{H}$. Thus the matroidal $\vec{w}$-fractional matching number $\tau_{w}(\mathcal{L})$ is equal to $\nu_{w}(\mathcal{H})$. And for any matroidal fractional matching $\vec{x}\in\cap_{i=1}^{k}P(\mathcal{M}_{i})=\cap_{i=1}^{k}Q(\mathcal{M}_{i})$, $x$ is a fractional matching of the hypergraph $\mathcal{H}$: for all the edges contains some $S_{i,j}$, the sum of their weights is just $x[S_{i,j}]$, which is at most $rank_{\mathcal{M}_{i}}(S_{i,j})=1$. Thus $\nu^{*}_{w}(\mathcal{L})\leq\nu_{w}^{*}(\mathcal{H})$.
Combined with Theorem 12.8 and Theorem 14.6, this argument yields
$\tau_{w}^{*}(\mathcal{L})=\nu_{w}^{*}(\mathcal{L})\leq\nu_{w}^{*}(\mathcal{H})\leq(k-1)\nu_{w}(\mathcal{H})=(k-1)\nu_{w}(\mathcal{L}),$
which completes the proof. ∎
By Corollary 9.7 and Theorem 14.1, the following two conjectures are equivalent to Conjecture 12.10.
Conjecture 14.8.
For any $\mathcal{L}\in\mathcal{D}^{k}$,
$R(\mathcal{L}):P(\bigcap\mathcal{L})\leq k-1.$
Conjecture 14.9.
For any =$\mathcal{L}=\{\mathcal{M}_{1},\cdots,\mathcal{M}_{k}\}\in\mathcal{D}^{k}$ on $V$$\mathcal{C}=\bigcap\mathcal{L}$, and $\vec{h}\in\mathbb{R}^{V}_{\geq 0}$,
$\chi^{*}(\mathcal{C},\vec{h})\leq(k-1)\max_{1\leq i\leq k}\chi^{*}(\mathcal{M}_{i},\vec{h}).$
Example 14.10 (Showing sharpness of  Conjecture 12.10, Conjecture 14.8, and Conjecture 14.9).
The truncated projective plane $T_{k}$ of uniformity $k$ is the induced subhypergraph $P_{k}[V(P_{k})\setminus\{v\}]$ for any vertex $v$ of a projective plane $P_{k}$ (assuming it exists). It is well-known that $T_{k}$ is $k$-partite and each part is of size $k-1$, each vertex is contained in $k-1$ edges, $\nu(T_{k})=1$, and $\nu^{*}(T_{k})=\tau(T_{k})=k-1$.
Let $\mathcal{L}(T_{k})\in\mathcal{D}^{k}$ be the system of partition matroids associated with $T_{k}$ and $\mathcal{C}=\bigcap\mathcal{L}(T_{k})$. Then $P(\mathcal{C})=Q(\mathcal{C})=\{x\in\mathbb{R}^{T_{k}}_{\geq 0}\mid x[T_{k}]\leq 1\}$. On the other hand, $\vec{y}=\frac{1}{k-1}\in Q(\mathcal{M}_{i})$ for each $1\leq i\leq k$, therefore $y\in R(\mathcal{L})$. But $y[T_{k}]=\frac{1}{k-1}(k-1)^{2}=k-1$. Thus $R(\mathcal{L}):P(\mathcal{C})\geq k-1$.
15. Fractional list colorings of complexes
In [8] a fractional version of list coloring of graphs was introduced. Here we generalize it to complexes.
Definition 15.1.
[…]
Clearly, $(\chi(\mathcal{C}),1)\in CL(\mathcal{C})$ so that $clr(\mathcal{C})\leq\chi(\mathcal{C})$.
Definition 15.2.
[…]
Clearly, $(\chi_{\ell}(\mathcal{C}),1)\in CH(\mathcal{C})$ so that $chr(\mathcal{C})\leq\chi_{\ell}(\mathcal{C})$.
Theorem 15.3.
For any complex $\mathcal{C}$, $\chi^{*}(\mathcal{C})=clr(\mathcal{C})=chr(\mathcal{C})$.
The proof is similar to that in  [8]. Details are given in Appendix C.
Appendix A A topological proof of Theorem 14.5
The proof below was given by Ori Kfir [20].
Remark A.1.
The proof yields the result also for “bipartite” hypergraphs, namely those having a set that meets every edge at one vertex. This is not an innovation, since Füredi’s original theorem contains also this case, because the condition it needs is non-containment of a projective plane, which obviously holds for “bipartite” hypergraphs.
A standard argument of adding dummy elements yields a deficiency version of Theorem 4.11:
Theorem A.2.
Let $V_{1},\ldots,V_{m}$ sets of vertices of a complex $\mathcal{C}$. If $\eta(\mathcal{C}[\cup_{i\in I}V_{i}])\geq|I|-d$ for every $I\subseteq[m]$ then there exists a partial transversal of size $m-d$, i.e., a choice function $\phi$ from $S\subseteq[m]$ of size $(m-d)$ to $\bigcup_{i\in S}F_{i}$ , such that $\phi(i)\in F_{i}$ for all $i\in S$ and $Im(\phi)\in\mathcal{C}$.
Theorem A.3 (Theorem 7.8 in [20]).
Let $\mathcal{H}$ be a $k$-uniform hypergraph. Suppose the vertex set of $\mathcal{H}$ can be divided into $U$ and $U^{\prime}$ such that for every edge $S\in\mathcal{H}$, $|S\cap U|=1$. Then
$\nu^{*}(\mathcal{H})\leq(k-1)\nu(\mathcal{H}).$
Proof.
Let $\mathcal{H}^{\prime}=\mathcal{H}/U=\{S\setminus U\mid S\in\mathcal{H}\}$, which is a $(k-1)$-uniform hypergraph by the assumption. For every $u\in U$ let $V_{u}=\{S\in\mathcal{H}^{\prime}\mid S\cup\{u\}\in\mathcal{H}\}$. Let
$d=\max_{W\subseteq U}\Big{(}|W|-\eta(\mathcal{M}[\cup_{w\in W}V_{w}])\Big{)},$
and let $Z$ be the subset of $W$ at which the maximum is attained. ($d\geq 0$ since $\eta(\emptyset)=0$.) By Theorem A.2, $\nu(\mathcal{H})\geq|U|-d$. On the other hand, by our choice, $\eta(\mathcal{M}[\cup_{w\in Z}V_{w}])=|Z|-d$. By Theorem 4.10
$(k-1)\eta(\mathcal{M}[\cup_{w\in Z}V_{w}])=(k-1)(|Z|-d)\geq\tau^{*}(\cup_{w\in Z}V_{Z}).$
Let $f$ be a fractional cover of $\cup_{w\in Z}V_{w}$ of total weight at most $(k-1)(|Z|-d)$. Then $f+\mathbf{1}_{U\setminus Z}$ is a fractional cover of $\mathcal{H}$, which has total weight at most
$(k-1)(|Z|-d)+|U|-|Z|\leq(k-1)(|Z|-d+|U|-|Z|)\leq(k-1)\nu(\mathcal{H}).$
Therefore $\nu^{*}(\mathcal{H})=\tau^{*}(\mathcal{H})\leq(k-1)\nu(\mathcal{H})$. ∎
Appendix B Remarks on Theorem 4.10
The result in [7] used a lower bound on $\eta$ formulated in terms of domination of independent sets. A larger lower bound is that of domination of fractional independent sets. There is no natural definition for those in general graphs: an attempt is “a function on the vertices summing up to no more than $1$ on every clique”, but this seems too liberal, so it does not produce any interesting results. But in the case of line graphs there is a natural definition, enabled by a tool called vector representations (of a graph). These are similar to Lovász  [23] function, but while the latter gives bounds on the independence number of a graph, ours is a vector representation of domination.
Definition B.1.
A function $P:V(G)\to\mathbb{R}^{n}$ is a vector representation of a graph $G$ if $P(u)\cdot P(v)\geq 1$ for every edge $\{u,v\}$ of $G$, and $P(G)\cdot P(v)\geq 0$ for every $u,v\in V(G)$.
A function $g:V(G)\rightarrow\mathbb{R}_{\geq 0}$ dominates $P$ if $\sum_{v\in V}g(v)P(v)\cdot P(u)\geq 1$ for every vertex $u\in V(G)$. The value of $P$ is defined as
$|P|=\min_{\text{$g$ dominates $P$}}g[V]$
and
$\Gamma(G)=\max_{\text{$P$ is a vector representation of~{}$G$}}|P|.$
In [4] the following was proved:
Theorem B.2.
For a graph $G$,
(35) $\eta(\mathcal{I}(G))\geq\Gamma(G).$
This inequality is particularly handy when it comes to line graphs, because a line graph has a canonical vector representation, in which every vertex $v$ is represented by the incidence vector of the edge to which it corresponds.
For a hypergraph $\mathcal{H}$, the fractional width $w^{*}(\mathcal{H})$ is the minimum of $\sum_{S\in\mathcal{H}}f(S)$ over all $f:\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ with the property that for every edge $S\in\mathcal{H}$,
$\sum_{T\in\mathcal{H}}f(T)\cdot|T\cap S|\geq 1.$
A function $g$ attaining the fractional width dominates the canonical vector representation, hence the line graph $L(\mathcal{H})$ of a hypergraph $\mathcal{H}$ satisfies
(36) $\Gamma\Big{(}L(\mathcal{H})\Big{)}\geq w^{*}(\mathcal{H}).$
On the other hand, the LP dual of fractional width of a hypergraph $\mathcal{H}$ is the maximum of $\sum_{S\in\mathcal{H}}g(S)$ over all $g:\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ with the property that for every edge $S\in\mathcal{H}$,
$\sum_{T\in\mathcal{H}}g(T)\cdot|T\cap S|\leq 1.$
Let $f$ be a fractional matching of $\mathcal{H}$ with $f[\mathcal{H}]=\nu^{*}(\mathcal{H})$. Suppose $\mathcal{H}$ is $k$-uniform. Set $g=\frac{f}{k}$. Then $\sum_{S\in\mathcal{H}}g(S)=\frac{\nu^{*}(\mathcal{H})}{k}$ and for every $S\in\mathcal{H}$,
$\sum_{T\in\mathcal{H}}g(T)\cdot|T\cap S|=\sum_{v\in S}\Big{(}\sum_{T\in\mathcal{H}:v\in T}g(T)\Big{)}\leq k\cdot\frac{1}{k}=1,$
where the first inequality is by the fractional matching property of $f$. Therefore
(37) $w^{*}(\mathcal{H})\geq\frac{\nu^{*}(\mathcal{H})}{k}.$
Combining (35), (36), and (37), it follows that for a $k$-uniform hypergraph $\mathcal{H}$,
$\eta\Big{(}\mathcal{M}(\mathcal{H})\Big{)}=\eta\Big{(}\mathcal{I}\big{(}L(\mathcal{H})\big{)}\Big{)}\geq\Gamma\Big{(}L(\mathcal{H})\Big{)}\geq w^{*}(\mathcal{H})\geq\frac{\nu^{*}(\mathcal{H})}{k},$
proving Theorem 4.10.
Appendix C Proof of Theorem 15.3
Lemma C.1.
For any complex $\mathcal{C}$, $clr(\mathcal{C})\geq\chi^{*}(\mathcal{C})$.
Proof.
For any $(a,b)\in CL(\mathcal{C})$, we have to show $\chi^{*}(\mathcal{C})\leq\frac{a}{b}$. By the $(a,b)$-colorable property, for the special $a$-list $L_{v}=\{1,2,\dots,a\}$ for each $v$, we can choose a $b$-set $C_{v}\subseteq L_{v}$ for each $v$ such that $S_{i}=\{v\in V(\mathcal{C})\mid i\in C_{v}\}$ is in $\mathcal{C}$ for each $i\in\{1,2\dots,a\}$. Putting weight $z_{i}/b$ on each of $S_{i}$, where $z_{i}$ is the number of occurrences of $S_{i}$ in the multiset $\{S_{1},\dots,S_{a}\}$, and putting weight zero on all the other sets of $\mathcal{C}$, we obtain a fractional coloring, as each $v$ is in $b$ many of $\{S_{1},\dots,S_{a}\}$, and the fractional coloring is of total weight $a/b$. Therefore $\chi^{*}(\mathcal{C})\leq a/b$. ∎
Lemma C.2.
For a complex $\mathcal{C}$, $chr(\mathcal{C})\geq clr(\mathcal{C})$.
Proof.
It is because $(a,b)\in CH(\mathcal{C})$ implies $(a,b)\in CL(\mathcal{C})$. ∎
Lemma C.3.
For a complex $\mathcal{C}$, $\chi^{*}(\mathcal{C})\geq chr(\mathcal{C})$.
Proof.
The first observation is that the optimum of the linear program of fractional coloring can be obtained by some basic feasible solution $f$, and since the constraints of fractional coloring are integral, by Cramer’s rule, there exists a subset $\mathcal{C}_{0}$ of $\mathcal{C}$ and a positive integer $q$ such that for each $T\in\mathcal{C}_{0}$, $f(T)=p_{T}/q$ for some positive integer $p_{T}$, and for each $T\in\mathcal{C}\setminus\mathcal{C}_{0}$, $f(T)=0$. And
$\chi^{*}(\mathcal{C})=f[\mathcal{C}]=f[\mathcal{C}_{0}]=\frac{p}{q},$
where $p:=\sum_{T\in\mathcal{C}_{0}}p_{T}$. By taking each $T\in\mathcal{C}_{0}$ $p_{T}$ times, we obtain a collection of faces $\{T_{1},\dots,T_{p}\}$ (not necessarily distinct) such that each vertex $v\in V(\mathcal{C})$ is in at least $q$ many of them:
|$|\{j:v\in T_{j}\}|=\sum_{T\in\mathcal{C}_{0}:v\in T}p_{T}=q\sum_{T\in\mathcal{C}:v\in T}f(T)\geq q,$
where the last inequality is by the constraint of the linear program.
Then to prove $\frac{p}{q}=\chi^{*}(\mathcal{C})\geq chr(\mathcal{C})$, we have to show that for every $\epsilon>0$ there exists $(a,b)\in CH(\mathcal{C})$ such that $\frac{a}{b}\leq(1+\epsilon)\frac{p}{q}$. Consider $a:=(1+\epsilon)pt$ and $b:=qt$, for $t$ sufficiently large, and we can assume without loss of generality that $a,b$ are integers. For any size $a$ lists $(L_{v}:v\in V(\mathcal{C}))$, we take a random partition $\cup_{v\in V(\mathcal{C})}L_{v}=Z_{1}\cup\dots Z_{p}$ such that each color is included in each $Z_{j}$ independently with probability $1/p$. Noting that $p$ and $|V(\mathcal{C})|$ are fixed for any given $\mathcal{C}$, it is followed by the well-known Chernoff’s bounds that with positive probability, for each $v\in V(\mathcal{C})$ and $j\in\{1,\dots,p\}$, $|L_{v}\cap Z_{j}|$ is concentrated around its expectation $|L_{v}|/p=(1+\epsilon)t$. Thus we can take an $t$-subset $C_{v,j}$ of $L_{v}\cap Z_{j}$. Then we get a set
$C_{v}:=\cup_{1\leq j\leq p:v\in T_{j}}C_{v,j}$
of size at least $qt=b$.
Then for each color $i\in\cup_{v\in V(\mathcal{C})}L_{v}$ and $S_{i}=\{v\in V(\mathcal{C})\mid i\in C_{v}\}$, if $i\not\in\cup_{v\in V(\mathcal{C})}C_{v}$, then $S_{i}=\emptyset$, otherwise $S_{i}\subseteq T_{j}$ for the index $j$ satisfying $i\in Z_{j}$. Since $T_{j}\in\mathcal{C}$, in either case $S_{i}\in\mathcal{C}$ and we prove $(a,b)\in CH(\mathcal{C})$. ∎
Combining the above lemmas, we have
$\chi^{*}(\mathcal{C})=chr(\mathcal{C})=clr(\mathcal{C}),$
which completes the proof of Theorem 15.3.